### v2

# Sup-norm bounds for Jacobi cusp forms

Abstract.
In this article, we give bounds for the natural invariant norm of cusp forms of real weight $k$ and character $\chi$ for any cofinite Fuchsian subgroup $\Gamma\subset\mathrm{SL}_{2}(\mathbb{R})$. Using the representation of Jacobi cusp forms of integral weight $k$ and index $m$ for the modular group $\Gamma_{0}=\mathrm{SL}_{2}(\mathbb{Z})$ as linear combinations of modular forms of weight $k-\frac{1}{2}$ for some congruence subgroup of $\Gamma_{0}$ (depending on $m$) and suitable Jacobi theta functions, we derive bounds for the natural invariant norm of these Jacobi cusp forms. More specifically, letting $J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$ denote the complex vector space of Jacobi cusp forms under consideration and $\|\cdot\|_{\mathrm{Pet}}$ the pointwise Petersson norm on $J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$, we prove that for given $\epsilon>0$, the bound
$\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|f(\tau,z)\|_{\mathrm{Pet}}=O_{\epsilon}\big{(}k^{\frac{3}{4}}m^{1+\epsilon}\big{)}$
holds for any $f\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$, which is normalized with respect to the Petersson inner product, where the implied constant depends only on the choice of $\epsilon>0$.
2010 Mathematics Subject Classification:
11F11, 11F12
1. Introduction
1.1. Background
In general, bounds for automorphic forms and for their Fourier coefficients represent an area of great interest in number theory. More specifically, we mention in this respect the results of [FJK16], where J. Friedman, J. Jorgenson, and J. Kramer obtained optimal sup-norm bounds on average for cusp forms of even weight $k$ for any cofinite Fuchsian subgroup $\Gamma\subset\mathrm{SL}_{2}(\mathbb{R})$. These bounds turn out to be uniform with respect to the subgroup $\Gamma$. Moreover, in [FJK19], effective versions for these sup-norm bounds are given. With regard to sup-norm bounds for individual Hecke eigenforms of large level, we mention, for example, the results by V. Blomer and R. Holowinsky in [BH11].
So far, less attention has been devoted to the study of sup-norm bounds for Jacobi cusp forms. The first comprehensive study of Jacobi forms was undertaken by M. Eichler and D. Zagier in [EZ85]. Subsequently, various authors have built on their work. In contrast to their analytical approach, a geometrical approach to the theory of Jacobi forms was given by J. Kramer in [Kr91].
Let $k,m$ be positive integers. A Jacobi form of weight $k$ and index $m$ for the modular group $\Gamma_{0}\coloneqq\mathrm{SL}_{2}(\mathbb{Z})$ is a holomorphic function on the product $\mathbb{H}\times\mathbb{C}$ of the upper half-plane $\mathbb{H}$ with the complex plane $\mathbb{C}$ having a suitable transformation behaviour with respect to $\Gamma_{0}$ and vanishing “at infinity”. We denote the complex vector space of Jacobi cusp forms of weight $k$ and index $m$ for $\Gamma_{0}$ by $J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$. The pointwise Petersson norm of a Jacobi form $f\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$ is then defined by
$\|f(\tau,z)\|^{2}_{\mathrm{Pet}}\coloneqq|f(\tau,z)|^{2}\,\mathrm{Im}(\tau)^{k}\,e^{-\frac{4\pi m\mathrm{Im}(z)^{2}}{\mathrm{Im}(\tau)}}\qquad(\tau\in\mathbb{H},\,z\in\mathbb{C}).$
Let $F$ be a Siegel cusp form of weight $k$ for the Siegel modular group $\mathrm{Sp}_{4}(\mathbb{Z})$, and let $\{f_{m}\}_{m\geq 1}$ be the set of Jacobi forms appearing in the Fourier–Jacobi expansion of $F$, i. e., $f_{m}\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$. Then, for any $\epsilon>0$, W. Kohnen proved the following sup-norm bound for the pointwise Petersson norm of $f_{m}$ in [Ko93]
(1) $\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|f_{m}(\tau,z)\|_{\mathrm{Pet}}=O_{F,\epsilon}\big{(}m^{\frac{k}{2}-\frac{2}{9}+\epsilon}\big{)},$
where the implied constant depends on the Siegel cusp form $F$ and on the choice of $\epsilon>0$. Motivated by the Ramanujan–Petersson conjecture, W. Kohnen then conjectured the bound
$\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|f_{m}(\tau,z)\|_{\mathrm{Pet}}=O_{F,\epsilon}\big{(}m^{\frac{k-1}{2}+\epsilon}\big{)},$
where the implied constant depends on the Siegel cusp form $F$ and on the choice of $\epsilon>0$.
More recently, P. Anamby and S. Das established in [AD23] a general sup-norm bound for the pointwise Petersson norm of $f\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$, which is normalized with respect to the Petersson inner product, i. e., for which we have
$\int\limits\limits_{\Gamma_{0}\ltimes\mathbb{Z}^{2}\backslash\mathbb{H}\times\mathbb{C}}\|f(\tau,z)\|^{2}_{\mathrm{Pet}}\,\frac{\mathrm{d}\xi\wedge\mathrm{d}\eta\wedge\mathrm{d}x\wedge\mathrm{d}y}{\eta^{3}}=1\qquad(\tau=\xi+i\eta,\,z=x+iy).$
Their bound is (see Theorem 1.4 in[AD23])
(2) $\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|f(\tau,z)\|_{\mathrm{Pet}}=O_{\epsilon}\big{(}k\,m\big{)},$
where the implied constant depends on the choice of $\epsilon>0$.
1.2. Main results
The goal of this article is to provide new sup-norm bounds for the pointwise Petersson norm for Jacobi forms of integral weight $k$ and integral index $m$ for $\Gamma_{0}$, which are normalized with respect to the Petersson inner product. The main result in this respect is given in Theorem 4.4 and states for $f\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$, which is normalized with respect to the Petersson inner product, that
(3) $\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|f(\tau,z)\|_{\mathrm{Pet}}=O_{\epsilon}\big{(}k^{\frac{3}{4}}\,m^{1+\epsilon}\big{)},$
where the implied constant depends only on the choice of $\epsilon>0$. For the proof, we essentially use the representation of the Jacobi cusp forms under consideration as linear combinations of modular forms of weight $k-\frac{1}{2}$ for some congruence subgroup of $\Gamma_{0}$ (depending on $m$) and suitable Jacobi theta functions; we then need to derive bounds for the pointwise Petersson norms of these functions to arrive at our result. Comparing our bound (3) with the bound (2) by P. Anamby and S. Das, we realize an improvement with regard to the polynomial growth in $k$ by a factor of $k^{\frac{1}{4}}$, while the polynomial growth of the bound (2) in $m$ is better by a factor of $m^{\epsilon}$ (for any $\epsilon>0$).
In order to be able to derive our bound (3), we need sup-norm bounds for the pointwise Petersson norm of cusp forms of positive real weight $k$ and character $\chi$ for any cofinite Fuchsian subgroup $\Gamma$ of $\mathrm{SL}_{2}(\mathbb{R})$. Such bounds have been derived for finite index subgroups of $\mathrm{SL}_{2}(\mathbb{Z})$ by Steiner in [St16]. Although these bounds are uniform with respect to the index of the finite index subgroups of $\mathrm{SL}_{2}(\mathbb{Z})$, there are mild restrictions for a direct application of the bounds derived in [St16] to our setting. On the other hand, such bounds could also be derived from [FJK16] with some extra work. However, we provide here new, alternative proofs for the results of [FJK16] applying to any positive real weight $k$ and any character $\chi$ by using the Bergman kernel for the modular curve associated to $\Gamma$.
More specifically, given $\Gamma\subset\mathrm{SL}_{2}(\mathbb{R})$ a Fuchsian subgroup, $k\in\mathbb{R}_{>0}$, and $\chi\colon\Gamma\rightarrow\mathbb{C}^{\times}$ a character, we let $S_{k,\chi}(\Gamma)$ denote the space of cusp forms of weight $k$ and character $\chi$ for $\Gamma$. Denoting by $d_{k}$ the dimension of $S_{k,\chi}(\Gamma)$ and letting $\{f_{1},\ldots,f_{d_{k}}\}$ be an orthonormal basis of $S_{k,\chi}(\Gamma)$ with respect to the Petersson inner product, the Bergman kernel associated to $S_{k,\chi}(\Gamma)$ is then defined by
$B_{k,\chi}(\tau,\tau^{\prime})\coloneqq\sum_{j=1}^{d_{k}}f_{j}(\tau)\overline{f_{j}(\tau^{\prime})};$
it is straightforward that this definition does not depend on the choice of an orthonormal basis of $S_{k,\chi}(\Gamma)$. The pointwise Petersson norm of the Bergman kernel is defined by
$\|B_{k,\chi}(\tau,\tau^{\prime})\|_{\mathrm{Pet}}=|B_{k,\chi}(\tau,\tau^{\prime})|\,(\mathrm{Im}(\tau)\mathrm{Im}(\tau^{\prime}))^{\frac{k}{2}},$
which gives on the diagonal
(4) $\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}=\sum\limits_{j=1}^{d_{k}}\|f_{j}(\tau)\|^{2}_{\mathrm{Pet}}.$
As a second main result of this article, we establish in Theorem 3.3, assuming that $k\in\mathbb{R}_{\geq 5}$, for $\Gamma$ being cocompact without elliptic elements the bound
$\sup_{z\in\mathbb{H}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}=O_{\Gamma}(k);$
moreover, for $\Gamma$ being cofinite, we give the bound
$\sup_{z\in\mathbb{H}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}=O_{\Gamma}\big{(}k^{\frac{3}{2}}\big{)},$
where the implied constants depend only on the Fuchsian subgroup $\Gamma$. Due to the relation (4), these results reprove the sup-norm bounds on average obtained in [FJK16], but now for any real weight $k\in\mathbb{R}_{\geq 5}$ and any character $\chi$. Based on these results, we are then also able to prove the uniformity of the above bounds with respect to the subgroup $\Gamma$ in Theorem 3.5.
1.3. Outline
Let us briefly outline the contents of this article. In the subsequent, second section we collect all the necessary prerequisites for the sequel of the paper. In particular, we introduce the definitions of cusp forms and Jacobi cusp forms together with their respective (pointwise) Petersson inner products. Furthermore, we define the Bergman kernel for modular curves and state its basic properties. We close the section by recalling asymptotics of Bergman kernels associated to tensor powers of holomorphic line bundles on compact complex Kähler manifolds due to [Bo96], which are crucial in the derivation of the bound (3).
The third section is devoted to the revisiting of the sup-norm bounds on average obtained in [FJK16], but now for any real weight $k\in\mathbb{R}_{\geq 5}$ and any character $\chi$. Here, the proofs of Theorem 3.3 and Theorem 3.5 are provided.
In the fourth section, the bound (3) is proven in Theorem 4.4. In addition to some straightforward bounds established for the relative $\mathrm{L}^{2}$-norm of classical Jacobi theta functions, the above mentioned asymptotics of Bergman kernels associated to tensor powers of holomorphic line bundles corresponding to these Jacobi theta functions are crucial in the derivation of the proof of Theorem 4.4.
2. Preliminaries
2.1. Hyperbolic metric
Let $\mathbb{H}\coloneqq\{\tau\in\mathbb{C}\,|\,\tau=\xi+i\eta,\,\eta>0\}$ be the upper half-plane. We denote by $\mathrm{d}s^{2}_{\mathrm{hyp}}(\tau)$ the line element and by $\mu_{\mathrm{hyp}}(\tau)$ the volume form corresponding to the hyperbolic metric on $\mathbb{H}$, which is compatible with the complex structure of $\mathbb{H}$ and has constant curvature equal to $-1$. Locally on $\mathbb{H}$, we have
$\mathrm{d}s^{2}_{\mathrm{hyp}}(\tau)=\frac{\mathrm{d}\xi^{2}+\mathrm{d}\eta^{2}}{\eta^{2}}\quad\textrm{and}\quad\mu_{\mathrm{hyp}}(\tau)=\frac{\mathrm{d}\xi\wedge\mathrm{d}\eta}{\eta^{2}}\,.$
For $\tau,\tau^{\prime}\in\mathbb{H}$, we let $\mathrm{dist}_{\mathrm{hyp}}(\tau,\tau^{\prime})$ denote the hyperbolic distance between these two points. For later purposes, it is useful to introduce the displacement function
(5) $\sigma(\tau,\tau^{\prime})\coloneqq\cosh^{2}\bigg{(}\frac{\mathrm{dist}_{\mathrm{hyp}}(\tau,\tau^{\prime})}{2}\bigg{)}=\frac{|\tau-\overline{\tau}^{\prime}|^{2}}{4\,\mathrm{Im}(\tau)\mathrm{Im}(\tau^{\prime})}\,.$
2.2. Quotient space
Let $\Gamma\subset\mathrm{SL}_{2}(\mathbb{R})$ be a Fuchsian subgroup acting by fractional linear transformations on $\mathbb{H}$. Let $X_{\Gamma}$ be the quotient space $\Γ\Gamma\backslash\mathbb{H}$ and $g_{\Gamma}$ the genus of $X_{\Gamma}$. In the sequel, we identify $X_{\Gamma}$ with a fundamental domain $\mathcal{F}_{\Gamma}\subset\mathbb{H}$ for the group $\Gamma$, which we assume to be closed and connected.
Denote by
$\mathcal{P}_{\Gamma}=\{p_{1},\ldots,p_{s}\}$
the set of cusps of $\mathcal{F}_{\Gamma}$. Let $\sigma_{\mathcal{P},j}\in\mathrm{SL}_{2}(\mathbb{R})$ be a scaling matrix of the cusp $p_{j}$, that is, $p_{j}=\sigma_{\mathcal{P},j}i\infty$ with stabilizer subgroup $\Gamma_{p_{j}}$ described as
(6) $\sigma_{\mathcal{P},j}^{-1}\Gamma_{p_{j}}\sigma_{\mathcal{P},j}=\begin{cases}\bigg{\langle}\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\bigg{\rangle},&\text{ if }-\mathrm{id}\not\in\Gamma,\\ \bigg{\langle}\pm\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\bigg{\rangle},&\text{ if }-\mathrm{id}\in\Gamma,\end{cases}\qquad(j=1,\ldots,s).$
For $Y>0$, we let $\mathcal{F}^{Y}_{j}\subset\mathcal{F}_{\Gamma}$ denote the neighborhood of the cusp $p_{j}$ characterized by
$\sigma_{\mathcal{P},j}^{-1}\mathcal{F}^{Y}_{j}=\{\tau=\xi+i\eta\in\mathbb{H}\,|\,-1/2\leq\xi\leq 1/2,\,\eta\geq Y\}\qquad(j=1,\ldots,s).$
With these notations, we define $\mathcal{F}_{Y}$ to be the closure of the complement of the union $\mathcal{F}^{Y}_{1}\cup\ldots\cup\mathcal{F}^{Y}_{s}$ in $\mathcal{F}_{\Gamma}$, i. e.,
(7) $\mathcal{F}_{Y}\coloneqq\mathrm{cl}\big{(}\mathcal{F}_{\Gamma}\setminus\big{(}\mathcal{F}^{Y}_{1}\cup\ldots\cup\mathcal{F}^{Y}_{s}\big{)}\big{)},$
which is compact; we note that $\mathcal{F}_{Y}=\mathcal{F}_{\Gamma}$, if $\Gamma$ is cocompact. We choose $0<m_{Y}<M_{Y}$ such that for all $\tau\in\mathcal{F}_{Y}$ the inequalities
$m_{Y}\leq\mathrm{Im}(\sigma_{\mathcal{P},j}^{-1}\tau)\leq M_{Y}$
hold for all $j=1,\ldots,s$; we note that $m_{Y}$ and $M_{Y}$ depend on the choice of $Y$.
Denote by
$\mathcal{E}_{\Gamma}=\{e_{1},\ldots,e_{t}\}$
the set of elliptic fixed points of $\mathcal{F}_{\Gamma}$. Let $\Gamma_{e_{j}}$ and $m_{j}$ denote the stabilizer subgroup and order of the elliptic fixed point $e_{j}$, respectively.
We denote the hyperbolic length of the shortest closed geodesic on $X_{\Gamma}$ by $\ell_{\Gamma}$. For a domain $D\subset\mathbb{H}$, we denote its hyperbolic diameter by $\mathrm{diam}_{\mathrm{hyp}}(D)$ and its hyperbolic volume by $\mathrm{vol}_{\mathrm{hyp}}(D)$. Finally, the injectivity radius $r_{\Gamma}$ is defined by
(8) $r_{\Gamma}\coloneqq\inf\bigg{\{}\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)\,\bigg{|}\,\tau\in\mathcal{F}_{\Gamma},\,\gamma\in\Gamma\setminus\bigg{(}\bigcup_{j=1}^{s}\Gamma_{p_{j}}\cup\bigcup_{j=1}^{t}\Gamma_{e_{j}}\bigg{)}\bigg{\}}.$
We note that if $X_{\Gamma}$ is compact without elliptic fixed points, i. e., $\mathcal{P}_{\Gamma}=\mathcal{E}_{\Gamma}=\emptyset$, then the injectivity radius $r_{\Gamma}$ equals the length of the shortest closed geodesic $\ell_{\Gamma}$ of $X_{\Gamma}$.
2.3. Cusp forms and Bergman kernel
For $k\in\mathbb{R}_{>0}$ and a character $\chi\colon\Gamma\rightarrow\mathbb{C}^{\times}$, we let $S_{k,\chi}(\Gamma)$ denote the space of cusp forms of weight $k$ and character $\chi$ for $\Gamma$, i. e., the space of holomorphic functions $f\colon\mathbb{H}\rightarrow\mathbb{C}$, which have the transformation behavior
$f(\gamma z)(cz+d)^{-k}=\chi(\gamma)f(z)$
for all $\gamma=\big{(}\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\big{)}\in\Gamma$, and which vanish at all the cusps of $\mathcal{F}_{\Gamma}$. Given $f\in S_{k,\chi}(\Gamma)$, we define
$\|f(\tau)\|_{\mathrm{Pet}}^{2}\coloneqq|f(\tau)|^{2}\,\eta^{k}\qquad(\tau=\xi+i\eta),$
which defines a $\Gamma$-invariant function on $\mathbb{H}$ called the pointwise Petersson norm of $f$.
The space $S_{k,\chi}(\Gamma)$ is equipped with the Petersson inner product
(9) $\langle f_{1},f_{2}\rangle_{\mathrm{Pet}}\coloneqq\int\limits_{\mathcal{F}_{\Gamma}}f_{1}(\tau)\overline{f_{2}(\tau)}\,\eta^{k}\mu_{\mathrm{hyp}}(\tau)\qquad(f_{1},f_{2}\in S_{k,\chi}(\Gamma)).$
Let $d_{k}$ denote the dimension of $S_{k,\chi}(\Gamma)$ and let $\{f_{1},\ldots,f_{d_{k}}\}$ be an orthonormal basis of $S_{k,\chi}(\Gamma)$ with respect to the Petersson inner product. Then, the Bergman kernel associated to $S_{k,\chi}(\Gamma)$ is defined by
$B_{k,\chi}(\tau,\tau^{\prime})\coloneqq\sum_{j=1}^{d_{k}}f_{j}(\tau)\overline{f_{j}(\tau^{\prime})}.$
It is obvious that this definition does not depend on the choice of an orthonormal basis of $S_{k,\chi}(\Gamma)$.
The Bergman kernel $B_{k,\chi}(\tau,\tau^{\prime})$ is a holomorphic cusp form of weight $k$ and character $\chi$ for $\Gamma$ in the $\tau$-variable, and an anti-holomorphic cusp form of weight $k$ and character $\overline{\chi}$ for $\Gamma$ in the $\tau^{\prime}$-variable. Hence, the pointwise Petersson norm of the Bergman kernel is given by
$\|B_{k,\chi}(\tau,\tau^{\prime})\|_{\mathrm{Pet}}=|B_{k,\chi}(\tau,\tau^{\prime})|\,(\eta\eta^{\prime})^{\frac{k}{2}},$
which is a $\Gamma$-invariant function on $\mathbb{H}\times\mathbb{H}$ with respect to both variables.
Moreover, $B_{k,\chi}(\tau,\tau^{\prime})$ is the reproducing kernel for $S_{k,\chi}(\Gamma)$, i. e., we have
$\int\limits_{\mathcal{F}_{\Gamma}}B_{k,\chi}(\tau,\tau^{\prime})f(\tau^{\prime})\eta^{\prime k}\mu_{\mathrm{hyp}}(\tau^{\prime})=f(\tau)\qquad(\tau^{\prime}=\xi^{\prime}+i\eta^{\prime})$
for any $f\in S_{k,\chi}(\Gamma)$. Therefore, for $k\in\mathbb{R}_{>3}$, the Bergman kernel $B_{k,\chi}(\tau,\tau^{\prime})$ can also be represented in the following form (see Proposition 1.3 on p. 77 in [Fr90])
(10) $B_{k,\chi}(\tau,\tau^{\prime})=\frac{(2i)^{k}(k-1)}{4\pi}\sum_{\gamma=\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)\in\Gamma}\frac{1}{(\tau-\gamma\overline{\tau}^{\prime})^{k}}\frac{1}{\chi(\gamma)(c\overline{\tau}^{\prime}+d)^{k}}\,.$
Note that the formula for the Bergman kernel given in [Fr90] is missing a factor of $(2i)^{k}$.
2.4. Counting function
Given $\tau\in\mathbb{H}$ and $\rho\in\mathbb{R}_{\geq 0}$, we recall from [JL95] the counting function
$N_{\Gamma}(\tau;\rho)\coloneqq|\mathcal{N}_{\Gamma}(\tau;\rho)|,$
where
$\mathcal{N}_{\Gamma}(\tau;\rho)\coloneqq\bigg{\{}\gamma\in\Gamma\setminus\bigg{(}\bigcup_{j=1}^{s}\Gamma_{p_{j}}\cup\bigcup_{j=1}^{t}\Gamma_{e_{j}}\bigg{)}\,\bigg{|}\,\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)\leq\rho\bigg{\}}.$
Let now $f$ be a positive, smooth, and decreasing function on $\mathbb{R}_{\geq 0}$. Then, adapting the arguments from [JL95] to Fuchsian subgroups of $\mathrm{SL}_{2}(\mathbb{R})$, we have for any $\tau\in\mathbb{H}$ and any $\delta\geq r_{\Gamma}/2$ the inequality
(11) $\int\limits_{0}^{\infty}f(\rho)\,\mathrm{d}N_{\Gamma}(\tau;\rho)\leq$ $\int\limits_{0}^{\delta}f(\rho)\,\mathrm{d}N_{\Gamma}(\tau;\rho)+\frac{2|\mathrm{Cent}(\Gamma)|\cosh(r_{\Gamma}/4)}{\sinh(r_{\Gamma}/4)}\sinh(\delta)f(\delta)$
$+\frac{|\mathrm{Cent}(\Gamma)|}{2\sinh^{2}(r_{\Gamma}/4)}\int\limits_{\delta}^{\infty}f(\rho)\sinh(\rho+r_{\Gamma}/2\big{)}\,\mathrm{d}\rho;$
here $\mathrm{Cent}(\Gamma)$ denotes the center of $\Gamma$. Note that our definition (8) of injectivity radius differs from the one used in [JL95] by a factor of $2$, and the inequality (11) takes this fact into account.
2.5. Jacobi forms
For $k,m\in\mathbb{N}$, we let $J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$ denote the space of Jacobi cusp forms of weight $k$ and index $m$ for $\Gamma_{0}=\mathrm{SL}_{2}(\mathbb{Z})$, i. e., the space of holomorphic functions $f\colon\mathbb{H}\times\mathbb{C}\rightarrow\mathbb{C}$, which have the transformation behaviour
$f\Big{(}\frac{a\tau+b}{c\tau+d},\frac{z+\lambda\tau+\mu}{c\tau+d}\Big{)}(c\tau+d)^{-k}\exp\Big{(}2\pi im\Big{(}\lambda^{2}\tau+2\lambda z-\frac{c(z+\lambda\tau+\mu)^{2}}{c\tau+d}\Big{)}\Big{)}=f(\tau,z)$
for all $\big{[}\big{(}\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\big{)},(\lambda,\mu)\big{]}\in\Gamma_{0}\ltimes\mathbb{Z}^{2}$, and which have a Fourier expansion of the form
$f(\tau,z)=\sum\limits_{\begin{subarray}{c}n\in\mathbb{N},\,r\in\mathbb{Z}\\ 4mn-r^{2}>0\end{subarray}}c(n,r)\,q^{n}\zeta^{r}\qquad(q=e^{2\pi i\tau},\,\zeta=e^{2\pi iz}).$
Given $f\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$, we define
$\|f(\tau,z)\|_{\mathrm{Pet}}^{2}\coloneqq|f(\tau,z)|^{2}\,\eta^{k}\,e^{-\frac{4\pi my^{2}}{\eta}}\qquad(\tau=\xi+i\eta,\,z=x+iy),$
which defines a $\Gamma_{0}\ltimes\mathbb{Z}^{2}$-invariant function on $\mathbb{H}\times\mathbb{C}$ called the pointwise Petersson norm of $f$.
Let $\mathcal{D}_{\Gamma_{0}}$ denote a fundamental domain of the quotient space $Y_{\Gamma_{0}}\coloneqq\Gamma_{0}\ltimes\mathbb{Z}^{2}\backslash\mathbb{H}\times\mathbb{C}$, which is a $2$-dimensional complex manifold. The space $J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$ is equipped with the Petersson inner product
(12) $\langle f_{1},f_{2}\rangle_{\mathrm{Pet}}\coloneqq\int\limits\limits_{\mathcal{D}_{\Gamma_{0}}}f_{1}(\tau,z)\overline{f_{2}(\tau,z)}\,\eta^{k}\,e^{-\frac{4\pi my^{2}}{\eta}}\,\frac{\mathrm{d}\xi\wedge\mathrm{d}\eta\wedge\mathrm{d}x\wedge\mathrm{d}y}{\eta^{3}}\qquad(f_{1},f_{2}\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})).$
For $f\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$, one has the decomposition
(13) $f(\tau,z)=\sum_{\mu=0}^{2m-1}\varphi_{\mu}(\tau)\vartheta_{\mu,m}(\tau,z),$
where the function $\varphi_{\mu}$ is a cusp form of weight $(k-\frac{1}{2})$ for the finite index subgroup $\Gamma_{1}\coloneqq\Gamma_{0}(4m)$ of $\Gamma_{0}$, and $\vartheta_{\mu,m}$ is the Jacobi theta function
(14) $\vartheta_{\mu,m}(\tau,z)\coloneqq\sum_{n\in\mathbb{Z}}e^{2\pi im\tau\big{(}n-\frac{\mu}{2m}\big{)}^{2}+2\pi iz(2mn-\mu)}.$
As we will see below, the theta functions $\vartheta_{\mu,m}$ ($\mu=0,\ldots,2m-1$) arise for fixed $\tau\in\mathbb{H}$ as global sections of a suitable line bundle on the elliptic curve associated to $\tau$. In fact, it is shown in Theorem 5.1 of [EZ85] that the decomposition (14) gives rise to the isomorphism
$J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})\cong\mathcal{V}_{k-\frac{1}{2}}(\Gamma_{0}),$
where $\mathcal{V}_{k-\frac{1}{2}}(\Gamma_{0})$ denotes the complex vector space of vector-valued cusp forms of weight $(k-\frac{1}{2})$ with suitable transformation behaviour with respect to $\Gamma_{0}$.
Let now
$f_{1}(\tau,z)=\sum_{\mu=0}^{2m-1}\varphi_{\mu,1}(\tau)\vartheta_{\mu,m}(\tau,z)\quad\text{and}\quad f_{2}(\tau,z)=\sum_{\mu=0}^{2m-1}\varphi_{\mu,2}(\tau)\vartheta_{\mu,m}(\tau,z)$
be two Jacobi cusp forms of weight $k$ and index $m$ for $\Gamma_{0}$. Then, the decomposition (13) gives rise to the equality (see Theorem 5.3 in [EZ85])
(15) $\langle f_{1},f_{2}\rangle_{\mathrm{Pet}}=\frac{1}{\sqrt{4m}}\int\limits_{\mathcal{F}_{\Gamma_{0}}}\sum_{\mu=0}^{2m-1}\varphi_{\mu,1}(\tau)\overline{\varphi_{\mu,2}(\tau)}\,\eta^{k-\frac{1}{2}}\,\frac{\mathrm{d}\xi\wedge\mathrm{d}\eta}{\eta^{2}},$
where we recall that $\mathcal{F}_{\Gamma_{0}}$ is a fundamental domain for the quotient space $X_{\Gamma_{0}}=\Gamma_{0}\backslash\mathbb{H}$.
2.6. Asymptotics for Bergman kernels
In this subsection, we recall asymptotics of Bergman kernels associated to tensor powers of holomorphic line bundles on compact complex Kähler manifolds, which are used in section 4 to derive bounds for theta functions.
Let $(M,\omega)$ be a compact complex Kähler manifold of dimension $n$ with positive closed $(1,1)$-form $\omega$. Let $\mathcal{L}$ be a positive hermitian holomorphic line bundle on $M$ and let $H^{0}(M,\mathcal{L}^{\otimes{m}})$ denote the vector space of global holomorphic sections its $m$-th tensor power $\mathcal{L}^{\otimes{m}}$ for $m\in\mathbb{Z}_{\geq 1}$. Let $|\,\cdot\,|_{\mathcal{L}^{\otimes{m}}}$ and $\langle\,\cdot\,,\,\cdot\,\rangle_{\mathrm{L}^{2},\,\mathcal{L}^{\otimes{m}}}$ denote the pointwise hermitian metric and the $\mathrm{L}^{2}$-inner product on $H^{0}(M,\mathcal{L}^{\otimes{m}})$, respectively.
Let $\{s_{j}\}$ denote an orthonormal basis of $H^{0}(M,\mathcal{L}^{\otimes{m}})$ with respect to the $\mathrm{L}^{2}$-inner product. For any $z\in M$, the function
$B_{\mathcal{L}^{\otimes{m}}}(z)\coloneqq\sum_{j}|s_{j}(z)|_{\mathcal{L}^{\otimes{m}}}^{2}$
is called the Bergman kernel associated to the line bundle $\mathcal{L}^{\otimes{m}}$. We note that the above definition is independent on the choice of an orthonormal basis of $H^{0}(M,\mathcal{L}^{\otimes{m}})$.
Let
(16) $\mathrm{c}_{1}(\mathcal{L},|\,\cdot\,|_{\mathcal{L}})(z)\coloneqq-\frac{i}{2\pi}\partial_{z}\partial_{\overline{z}}\log|s(z)|_{\mathcal{L}}^{2}$
denote the curvature form of the line bundle $\mathcal{L}$ at the point $z\in M$, where $s$ is any meromorphic section of $\mathcal{L}$. At any $z\in M$, there exists a coordinate chart around the point $z$ such that
$\omega(z)=\sum_{j=1}^{n}\frac{i}{2}\cdot\mathrm{d}z_{j}\wedge\mathrm{d}\overline{z}_{j}\quad\text{and}\quad\mathrm{c}_{1}(\mathcal{L},|\,\cdot\,|_{\mathcal{L}})(z)=\sum_{j=1}^{n}\frac{i}{2}\cdot\alpha_{j}\cdot\mathrm{d}z_{j}\wedge\mathrm{d}\overline{z}_{j}.$
The complex numbers $\alpha_{1},\ldots,\alpha_{n}$ are called the eigenvalues of the curvature form $\mathrm{c}_{1}(\mathcal{L},|\,\cdot\,|_{\mathcal{L}})(z)$ at the point $z\in M$. We set
$\mathrm{det}_{\omega}\big{(}\mathrm{c}_{1}(\mathcal{L},|\,\cdot\,|_{\mathcal{L}})(z)\big{)}\coloneqq\prod_{j=1}^{n}\alpha_{j}.$
Since the line bundle $\mathcal{L}$ is positive, we have $\alpha_{j}>0$ for $j=1,\ldots,n$. Finally, we recall from Theorem 2.1 in [Bo96] the bound
(17) $B_{\mathcal{L}^{\otimes{m}}}(z)=\mathrm{det}_{\omega}\big{(}\mathrm{c}_{1}(\mathcal{L},|\,\cdot\,|_{\mathcal{L}})(z)\big{)}m^{n}+O(m^{n-1}),$
provided that $\mathcal{L}$ is a positive line bundle for any $z\in M$.
3. Sup-norm bounds for cusp forms revisited
Refining arguments of [AM17] and [AM18], we first derive bounds for the Bergman kernel along the diagonal.
Proposition 3.1.
With notations as above, let $\Gamma\subset\mathrm{SL}_{2}(\mathbb{R})$ be a cocompact Fuchsian subgroup without elliptic elements. Then, for $k\in\mathbb{R}_{\geq 5}$ and $\tau\in\mathbb{H}$, we have the bound
$\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}\leq\frac{k-1}{2\pi}+\frac{3(k-1)}{\pi\cosh^{k-4}(r_{\Gamma}/4)}\bigg{(}1+\frac{1}{\sinh^{2}(r_{\Gamma}/4)}\bigg{)}.$
Proof.
Letting $k\in\mathbb{R}_{\geq 5}$ and considering the Bergman kernel (10) on the diagonal, we derive by means of relation (5) the bound
$\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}=\frac{2^{k}(k-1)}{4\pi}\bigg{|}\sum_{\gamma=\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)\in\Gamma}\frac{1}{(\tau-\gamma\overline{\tau})^{k}}\frac{1}{\chi(\gamma)(c\overline{\tau}+d)^{k}}\bigg{|}\,\mathrm{Im}(\tau)^{k}$
$\qquad\leq\frac{k-1}{4\pi}\sum_{\gamma\in\Gamma}\bigg{(}\frac{4\,\mathrm{Im}(\tau)\mathrm{Im}(\gamma\tau)}{|\tau-\gamma\overline{\tau}|^{2}}\bigg{)}^{k/2}=\frac{k-1}{4\pi}\sum_{\gamma\in\Gamma}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}$
(18) $\qquad=\frac{k-1}{4\pi}\bigg{(}|\mathrm{Cent}(\Gamma)|+\sum_{\gamma\in\Gamma\setminus\mathrm{Cent}(\Gamma)}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}\bigg{)}.$
Substituting $\delta=r_{\Gamma}/2$ in inequality (11) and using the fact that $|\mathrm{Cent}(\Gamma)|\leq 2$, we derive
$\sum_{\gamma\in\Gamma\setminus\mathrm{Cent}(\Gamma)}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}$
(19) $\qquad\leq\int\limits_{0}^{r_{\Gamma}/2}\frac{\mathrm{d}N_{\Gamma}(\tau;\rho)}{\cosh^{k}(\rho/2)}+\frac{8}{\cosh^{k-2}(r_{\Gamma}/4)}+\frac{1}{\sinh^{2}(r_{\Gamma}/4)}\int\limits_{r_{\Gamma}/2}^{\infty}\frac{\sinh(\rho+r_{\Gamma}/2)}{\cosh^{k}(\rho/2)}\,\mathrm{d}\rho.$
From the defining equation (8) of the injectivity radius $r_{\Gamma}$, we find for the first term of (19) that
(20) $\int\limits_{0}^{r_{\Gamma}/2}\frac{\mathrm{d}N_{\Gamma}(\tau;\rho)}{\cosh^{k}(\rho/2)}=0.$
With regard to the third term of (19), we recall the bound (12) in [AM17], which states for any $k\in\mathbb{R}_{\geq 5}$ and any $\delta\geq 0$ (note that we have replaced $2k$ by $k$) that
$\frac{1}{\sinh^{2}(r_{\Gamma}/4)}\int\limits_{\delta}^{\infty}\frac{\sinh(\rho+r_{\Gamma}/2)}{\cosh^{k}(\rho/2)}\,\mathrm{d}\rho$
(21) $\quad\leq\frac{4}{(k-2)\cosh^{k-2}(\delta/2)}\bigg{(}2+\frac{1}{\sinh^{2}(r_{\Gamma}/4)}\bigg{)}+\frac{8}{(k-4)\cosh^{k-4}(\delta/2)}\cdot\frac{1}{\sinh^{2}(r_{\Gamma}/4)}.$
From the elementary inequality $\cosh^{k-4}(r_{\Gamma}/4)\leq\cosh^{k-2}(r_{\Gamma}/4)$ and recalling that $k\in\mathbb{R}_{\geq 5}$, we now derive from (21) with $\delta=r_{\Gamma}/2$ the bound
$\frac{1}{\sinh^{2}(r_{\Gamma}/4)}\int\limits_{r_{\Gamma}/2}^{\infty}\frac{\sinh(\rho+r_{\Gamma}/2)}{\cosh^{k}(\rho/2)}\,\mathrm{d}\rho$
$\leq\frac{4}{(k-2)\cosh^{k-2}(r_{\Gamma}/4)}\bigg{(}2+\frac{1}{\sinh^{2}(r_{\Gamma}/4)}\bigg{)}+\frac{8}{(k-4)\cosh^{k-4}(r_{\Gamma}/4)}\cdot\frac{1}{\sinh^{2}(r_{\Gamma}/4)}$
(22) $\leq\frac{4}{\cosh^{k-4}(r_{\Gamma}/4)}\bigg{(}1+\frac{1}{\sinh^{2}(r_{\Gamma}/4)}\bigg{)}+\frac{8}{(k-4)\cosh^{k-4}(r_{\Gamma}/4)}\cdot\frac{1}{\sinh^{2}(r_{\Gamma}/4)}.$
Combining the bounds (18), (19) with (20), (22), and using the fact that $k\in\mathbb{R}_{\geq 5}$, completes the proof of the proposition. ∎
Proposition 3.2.
With notations as above, let $\Gamma\subset\mathrm{SL_{2}(\mathbb{R})}$ be a cofinite Fuchsian subgroup. Then, for $k\in\mathbb{R}_{\geq 5}$ and $\tau\in\mathbb{H}$, we have the bound
$\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}$ $\leq\frac{k-1}{2\pi}+\frac{3(k-1)}{\pi\cosh^{k-4}(r_{\Gamma}/4)}\bigg{(}1+\frac{1}{\sinh^{2}(r_{\Gamma}/4)}\bigg{)}$
$\quad+\frac{k-1}{4\pi}\sum_{e_{j}\in\mathcal{E}_{\Gamma}}(m_{j}-1)+\frac{2(k-1)}{\sqrt{\pi}}\cdot\frac{\Gamma((k-1)/2)}{\Gamma(k/2)}\sum_{p_{j}\in\mathcal{P}_{\Gamma}}\mathrm{Im}(\sigma_{\mathcal{P},j}^{-1}\tau),$
where $\sigma_{\mathcal{P},j}$ is the scaling matrix associated to the cusp $p_{j}\in\mathcal{P}_{\Gamma}$ defined in (6).
Proof.
For $k\in\mathbb{R}_{\geq 5}$ and $\tau\in\mathbb{H}$, using the bound (18) and the fact that $|\mathrm{Cent}(\Gamma)|\leq 2$, we derive
$\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}$ $\leq\frac{k-1}{2\pi}+\frac{k-1}{4\pi}\sum_{\gamma\in\Gamma\setminus\big{(}\bigcup_{e_{j}\in\mathcal{E}_{\Gamma}}\Gamma_{e_{j}}\,\cup\,\bigcup_{p_{j}\in\mathcal{P}_{\Gamma}}\Gamma_{p_{j}}\big{)}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}$
$\quad+\frac{k-1}{4\pi}\sum_{e_{j}\in\mathcal{E}_{\Gamma}}\,\sum_{\gamma\in\Gamma_{e_{j}}\setminus\mathrm{Cent}(\Gamma)}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}$
(23) $\quad+\frac{k-1}{4\pi}\sum_{p_{j}\in\mathcal{P}_{\Gamma}}\,\sum_{\gamma\in\Gamma_{p_{j}}\setminus\mathrm{Cent}(\Gamma)}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}.$
Adapting our arguments from Proposition 3.1 to the second summand on the right-hand side of (23), we arrive at the bound
$\frac{k-1}{4\pi}\sum_{\gamma\in\Gamma\setminus\big{(}\bigcup_{e_{j}\in\mathcal{E}_{\Gamma}}\Gamma_{e_{j}}\,\cup\,\bigcup_{p_{j}\in\mathcal{P}_{\Gamma}}\Gamma_{p_{j}}\big{)}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}$
(24) $\leq\frac{3(k-1)}{\pi\cosh^{k-4}(r_{\Gamma}/4)}\bigg{(}1+\frac{1}{\sinh^{2}(r_{\Gamma}/4)}\bigg{)}.$
For the third term on the right-hand side of (23), we trivially have the bound
(25) $\frac{k-1}{4\pi}\sum_{e_{j}\in\mathcal{E}_{\Gamma}}\,\sum_{\gamma\in\Gamma_{e_{j}}\setminus\mathrm{Cent}(\Gamma)}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}\leq\frac{k-1}{4\pi}\sum_{e_{j}\in\mathcal{E}_{\Gamma}}(m_{j}-1).$
From the definition of the scaling matrix (6) and using the fact that $|\mathrm{Cent(\Gamma)}|\leq 2$, we find
$\frac{k-1}{4\pi}\sum_{p_{j}\in\mathcal{P}_{\Gamma}}\,\sum_{\gamma\in\Gamma_{p_{j}}\setminus\mathrm{Cent}(\Gamma)}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}$
(26) $\qquad\leq\frac{k-1}{2\pi}\sum_{p_{j}\in\mathcal{P}_{\Gamma}}\,\sum_{n\in\mathbb{Z}\backslash\{0\}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\sigma_{\mathcal{P},j}^{-1}\tau,\sigma_{\mathcal{P},j}^{-1}\tau+n)/2)}.$
We now recall the bound (18) in [AM17], which gives for $k\in\mathbb{R}_{\geq 5}$, $p_{j}\in\mathcal{P}_{\Gamma}$, and $\tau,\tau^{\prime}\in\mathbb{H}$ (note that we have replaced $2k$ by $k$) the bound
$\frac{k-1}{2\pi}\sum_{p_{j}\in\mathcal{P}_{\Gamma}}\,\sum_{n\in\mathbb{Z}\backslash\{0\}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\sigma_{\mathcal{P},j}^{-1}\tau,\sigma_{\mathcal{P},j}^{-1}\tau^{\prime}+n)/2)}$
(27) $\qquad\leq\frac{k-1}{\sqrt{\pi}}\cdot\frac{\Gamma\big{(}(k-1)/2\big{)}}{\Gamma(k/2)}\sum_{p_{j}\in\mathcal{P}_{\Gamma}}\frac{\big{(}4\,\mathrm{Im}(\sigma_{\mathcal{P},j}^{-1}\tau)\mathrm{Im}(\sigma_{\mathcal{P},j}^{-1}\tau^{\prime})\big{)}^{k/2}}{\big{(}\mathrm{Im}(\sigma_{\mathcal{P},j}^{-1}\tau)+\mathrm{Im}(\sigma_{\mathcal{P},j}^{-1}\tau^{\prime})\big{)}^{k-1}}.$
Substituting $\tau=\tau^{\prime}$ in (27) and combining it with (26), we arrive for the fourth term on the right-hand side of (23) at the bound
$\frac{k-1}{4\pi}\sum_{p_{j}\in\mathcal{P}_{\Gamma}}\,\sum_{\gamma\in\Gamma_{p_{j}}\setminus\mathrm{Cent}(\Gamma)}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}$
(28) $\qquad\leq\frac{2(k-1)}{\sqrt{\pi}}\cdot\frac{\Gamma\big{(}(k-1)/2\big{)}}{\Gamma(k/2)}\sum_{p_{j}\in\mathcal{P}_{\Gamma}}\mathrm{Im}(\sigma_{\mathcal{P},j}^{-1}\tau).$
Combining the bounds (24), (25), and (28) with (23) completes the proof of the proposition. ∎
Theorem 3.3.
With notations as above, let $\Gamma\subset\mathrm{SL_{2}(\mathbb{R})}$ be a cofinite Fuchsian subgroup and $k\in\mathbb{R}_{\geq 5}$. Then, if $\Gamma$ is cocompact without elliptic elements, we have the bound
(29) $\sup_{\tau\in\mathbb{H}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}=O_{\Gamma}(k).$
Moreover, if $\Gamma$ is cofinite, we have the bound
(30) $\sup_{\tau\in\mathbb{H}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}=O_{\Gamma}\big{(}k^{\frac{3}{2}}\big{)}.$
The implied constants in the bounds (29) and (30) depend only on the Fuchsian subgroup $\Gamma$.
Proof.
When $\Gamma$ is cocompact without elliptic elements, the claimed bound (29) follows directly from Proposition 3.1.
Let next $\Gamma$ be a cofinite Fuchsian subgroup. From the proof of Theorem 6.1 in [FJK16], it follows that
(31) $\sup_{\tau\in\mathbb{H}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}=\sup_{\begin{subarray}{c}\tau\in\partial\mathcal{F}_{Y}\\ Y=k/(2\pi)\end{subarray}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}},$
where $\partial\mathcal{F}_{Y}$ denotes the boundary of the truncated fundamental domain $\mathcal{F}_{Y}$ defined in (7). Combining Proposition 3.2 with (31) and employing the fact that
$\frac{\Gamma\big{(}(k-1)/2\big{)}}{\Gamma(k/2)}=O\bigg{(}\frac{1}{\sqrt{k}}\bigg{)},$
we derive
$\sup_{\tau\in\mathbb{H}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}$
$\qquad\leq\frac{k-1}{2\pi}+\frac{3(k-1)}{\pi\cosh^{k-4}(r_{\Gamma}/4)}\bigg{(}1+\frac{1}{\sinh^{2}(r_{\Gamma}/4)}\bigg{)}+(k-1)\,C_{\Gamma,\mathrm{ell}}+k^{\frac{3}{2}}\,C_{\Gamma,\mathrm{par}}$
for some positive constants $C_{\Gamma,\mathrm{ell}},C_{\Gamma,\mathrm{par}}$, which depend only on the Fuchsian subgroup $\Gamma$. This completes the proof of the theorem. ∎
Corollary 3.4.
With notations as above, let $\Gamma\subset\mathrm{SL_{2}(\mathbb{R})}$ be a cofinite Fuchsian subgroup. For $k\in\mathbb{R}_{\geq 5}$, let $f\in S_{k,\chi}(\Gamma)$ be a cusp form, which is normalized with respect to the Petersson inner product (9). If $\Gamma$ is cocompact without elliptic elements, we have the bound
(32) $\sup_{\tau\in\mathbb{H}}\|f(\tau)\|_{\mathrm{Pet}}^{2}=O_{\Gamma}(k).$
Moreover, if $\Gamma$ is cofinite, we have the bound
(33) $\sup_{\tau\in\mathbb{H}}\|f(\tau)\|_{\mathrm{Pet}}^{2}=O_{\Gamma}\big{(}k^{\frac{3}{2}}\big{)}.$
The implied constants in the bounds (32) and (33) depend only on the Fuchsian subgroup $\Gamma$.
Proof.
Choose an orthonormal basis $\{f_{1}=f,\ldots,f_{d_{k}}\}$ of $S_{k,\chi}(\Gamma)$ with respect to the Petersson inner product (9). For $\tau\in\mathbb{H}$, we then have the bound
$\|f(\tau)\|_{\mathrm{Pet}}^{2}\leq\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}.$
The proof of the corollary now immediately follows from Theorem 3.3. ∎
Theorem 3.5.
With notations as above, let $\Gamma_{0}\subset\mathrm{SL_{2}}(\mathbb{R})$ be a fixed cofinite Fuchsian subgroup and let $\Gamma\subseteq\Gamma_{0}$ be a finite index subgroup of $\Gamma_{0}$. For $k\in\mathbb{R}_{\geq 5}$, let $f\in S_{k,\chi}(\Gamma)$ be a cusp form, which is normalized with respect to the Petersson inner product (9). If $\Gamma_{0}$ is cocompact without elliptic elements, we have the bound
(34) $\sup_{\tau\in\mathbb{H}}\|f(\tau)\|_{\mathrm{Pet}}^{2}=O_{\Gamma_{0}}(k).$
Moreover, if $\Gamma_{0}$ is cofinite, we have the bound
(35) $\sup_{\tau\in\mathbb{H}}\|f(\tau)\|_{\mathrm{Pet}}^{2}=O_{\Gamma_{0}}\big{(}k^{\frac{3}{2}}\big{)}.$
The implied constants in the bounds (34) and (35) depend only on the Fuchsian subgroup $\Gamma_{0}$.
Proof.
Choose an orthonormal basis $\{f_{1}=f,\ldots,f_{d_{k}}\}$ of $S_{k,\chi}(\Gamma)$ with respect to the Petersson inner product (9).
Let now $\Gamma_{0}$ be a cocompact Fuchsian subgroup without elliptic elements. From the proof of Proposition 3.1,we derive the bound
$\sup_{\tau\in\mathbb{H}}\|f(\tau)\|_{\mathrm{Pet}}^{2}$ $\leq\sup_{\tau\in\mathbb{H}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}$
$\leq\frac{k-1}{4\pi}\sup_{\tau\in\mathbb{H}}\sum_{\gamma\in\Gamma}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}$
$\leq\frac{k-1}{4\pi}\sup_{\tau\in\mathbb{H}}\sum_{\gamma\in\Gamma_{0}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}=O_{\Gamma_{0}}(k),$
which completes the proof of the theorem in the case that $\Gamma_{0}$ is cocompact without elliptic elements.
Let next $\Gamma_{0}$ be a cofinite Fuchsian subgroup. Given $Y>0$, we recall from (7) the fundamental domain decomposition
$\mathcal{F}_{\Gamma}=\mathcal{F}_{Y}\cup\big{(}\mathcal{F}^{Y}_{1}\cup\ldots\cup\mathcal{F}^{Y}_{s}\big{)},$
where $\mathcal{F}_{Y}$ is a compact subset of $\mathcal{F}_{\Gamma}$ and the $\mathcal{F}^{Y}_{j}$’s are neighborhoods of the cusps $p_{j}\in\mathcal{P}_{\Gamma}$ ($j=1,\ldots,s$). Choosing $Y$ large enough, we can assume without loss of generality in the sequel that the neighborhoods $\mathcal{F}^{Y}_{j}$ are pairwise disjoint. We now first provide a bound for the pointwise Petersson norm of $f$, when $\tau$ ranges across the compact set $\mathcal{F}_{Y}$, and subsequently we compute bounds for the pointwise Petersson norm of $f$, when $\tau$ ranges across the neighborhoods $\mathcal{F}^{Y}_{j}$ of the cusps for fixed, large enough $Y$.
Adapting arguments from the proof of Proposition 3.2, we obtain the bound
$\sup_{\tau\in\mathcal{F}_{Y}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}\leq\frac{k-1}{4\pi}\sup_{\tau\in\mathcal{F}_{Y}}\sum_{\gamma\in\Gamma}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}$
$\qquad\,\leq\frac{k-1}{4\pi}\sup_{\tau\in\mathcal{F}_{Y}}\sum_{\gamma\in\Gamma_{0}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}$
(36) $\qquad\leq\frac{k-1}{2\pi}+\frac{3(k-1)}{\pi\cosh^{k-4}(r_{\Gamma_{0},Y}/4)}\bigg{(}1+\frac{1}{\sinh^{2}(r_{\Gamma_{0},Y}/4)}\bigg{)}+\frac{k-1}{4\pi}\sum_{e_{j}\in\mathcal{E}_{\Gamma_{0}}}(m_{e_{j}}-1),$
where
$r_{\Gamma_{0},Y}=\inf\bigg{\{}\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)\,\bigg{|}\,\tau\in\mathcal{F}_{Y},\,\gamma\in\Gamma_{0}\setminus\bigcup_{e_{j}\in\mathcal{E}_{\Gamma_{0}}}\Gamma_{0,e_{j}}\bigg{\}}>0.$
From this, we immediately conclude that
(37) $\sup_{\tau\in\mathcal{F}_{Y}}\|f(\tau)\|_{\mathrm{Pet}}^{2}\leq\sup_{\tau\in\mathcal{F}_{Y}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}=O_{\Gamma_{0},Y}(k),$
where the implied constant depends on $\Gamma_{0}$ and the choice of $Y$.
We are left to provide bounds for the pointwise Petersson norm of $f$, when $\tau$ ranges across the neighborhoods $\mathcal{F}^{Y}_{j}$ of the cusps. For this, we will have to distinguish between the two cases $Y>\frac{k}{2\pi}$ and $Y<\frac{k}{2\pi}$. Without loss of generality, we can assume that $j=1$, when $p_{1}\in\mathcal{P}_{\Gamma}$ is the cusp at infinity for $\Gamma$ lying above the cusp $p$ at infinity for $\Gamma_{0}$ with ramification index $[\Gamma_{0,p}:\Gamma_{p_{1}}]$ (note that $\Gamma_{0,p}$ denotes the stabilizer subgroup of $p$ in $\Gamma_{0}$). With the above notations, we obtain the inclusion $\Gamma\setminus\Gamma_{p_{1}}\subseteq\Gamma_{0}\setminus\Gamma_{0,p}$.
We first treat the case $Y>\frac{k}{2\pi}$, which implies that $\mathcal{F}_{1}^{Y}\subset\mathcal{F}_{1}^{k/(2\pi)}$. Arguing as in the proof of Theorem 6.1 in [FJK16], we deduce, recalling the inclusion $\Gamma\setminus\Gamma_{p_{1}}\subseteq\Gamma_{0}\setminus\Gamma_{0,p}$, that
$\quad\sup_{\tau\in\mathcal{F}_{1}^{Y}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}\leq\sup_{\tau\in\mathcal{F}_{1}^{k/(2\pi)}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}\leq\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(2\pi)}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}$
$\leq\frac{k-1}{4\pi}\Bigg{(}\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(2\pi)}}\sum_{\gamma\in\Gamma\setminus\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}+\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(2\pi)}}\sum_{\gamma\in\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}\Bigg{)}$
$\leq\frac{k-1}{4\pi}\Bigg{(}\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(2\pi)}}\sum_{\gamma\in\Gamma_{0}\setminus\Gamma_{0,p}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}+\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(2\pi)}}\sum_{\gamma\in\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}\Bigg{)}.$
Arguments similar to the ones used to derive the bound (36), lead for the first term of (LABEL:thm6-eqn4) to the bound
$\frac{k-1}{4\pi}\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(2\pi)}}\sum_{\gamma\in\Gamma_{0}\setminus\Gamma_{0,p}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}$
$\quad\leq\frac{k-1}{2\pi}+\frac{3(k-1)}{\pi\cosh^{k-4}(r_{\Gamma_{0},k/(2\pi)}/4)}\bigg{(}1+\frac{1}{\sinh^{2}(r_{\Gamma_{0},k/(2\pi)}/4)}\bigg{)}+\frac{k-1}{4\pi}\sum_{e_{j}\in\mathcal{E}_{\Gamma_{0}}}(m_{e_{j}}-1),$
where
$r_{\Gamma_{0},k/(2\pi)}=\inf\bigg{\{}\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)\,\bigg{|}\,\tau\in\partial\mathcal{F}_{1}^{k/(2\pi)},\,\gamma\in\Gamma_{0}\setminus\bigg{(}\Gamma_{0,p}\cup\bigcup_{e_{j}\in\mathcal{E}_{\Gamma_{0}}}\Gamma_{0,e_{j}}\bigg{)}\bigg{\}}>0.$
Since it is easy to see that
$\frac{1}{\sinh^{2}(r_{\Gamma_{0},k/(2\pi)})}=O_{\Gamma_{0}}(1),$
we arrive for the first term of (LABEL:thm6-eqn4) at the bound
(39) $\frac{k-1}{4\pi}\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(2\pi)}}\sum_{\gamma\in\Gamma_{0}\setminus\Gamma_{0,p}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}=O_{\Gamma_{0}}(k).$
Using the same arguments as in the proof of Proposition 3.2, we derive for the second term of (LABEL:thm6-eqn4) the bound
(40) $\frac{k-1}{4\pi}\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(2\pi)}}\sum_{\gamma\in\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}=O\big{(}k^{\frac{3}{2}}\big{)}.$
By means of (LABEL:thm6-eqn4), we thus deduce from (39) and (40) in the case $Y>\frac{k}{2\pi}$ the bound
(41) $\sup_{\tau\in\mathcal{F}_{1}^{Y}}\|f(\tau)\|_{\mathrm{Pet}}^{2}\leq\sup_{\tau\in\mathcal{F}_{1}^{Y}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}=O_{\Gamma_{0}}\big{(}k^{\frac{3}{2}}\big{)}.$
We finally turn to the case $Y<\frac{k}{2\pi}$, which implies that $\mathcal{F}_{1}^{k/(2\pi)}\subset\mathcal{F}_{1}^{Y}$. Here we find, arguing as in the preceding case that
$\sup_{\tau\in\mathcal{F}_{1}^{Y}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}$ $\leq\sup_{\tau\in\mathcal{F}_{1}^{Y}\setminus\mathcal{F}_{1}^{k/(2\pi)}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}$
$\leq\frac{k-1}{4\pi}\sup_{\tau\in\mathcal{F}_{1}^{Y}\setminus\mathcal{F}_{1}^{k/(2\pi)}}\sum_{\gamma\in\Gamma_{0}\setminus\Gamma_{0,p}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}$
(42) $\quad\,+\frac{k-1}{4\pi}\sup_{\tau\in\mathcal{F}_{1}^{Y}\setminus\mathcal{F}_{1}^{k/(2\pi)}}\sum_{\gamma\in\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}.$
As before, we now obtain the bounds
$\frac{k-1}{4\pi}\sup_{\tau\in\mathcal{F}_{1}^{Y}\setminus\mathcal{F}_{1}^{k/(2\pi)}}\sum_{\gamma\in\Gamma_{0}\setminus\Gamma_{0,p}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}$
$\quad\leq\frac{k-1}{2\pi}+\frac{3(k-1)}{\pi\cosh^{k-4}(r^{\prime}_{\Gamma_{0},Y}/4)}\bigg{(}1+\frac{1}{\sinh^{2}(r^{\prime}_{\Gamma_{0},Y}/4)}\bigg{)}+\frac{k-1}{4\pi}\sum_{e_{j}\in\mathcal{E}_{\Gamma_{0}}}(m_{e_{j}}-1)$
(43) $\quad=O_{\Gamma_{0},Y}(k),$
noting that
$r^{\prime}_{\Gamma_{0},Y}=\inf\bigg{\{}\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)\,\bigg{|}\,\tau\in\mathcal{F}_{1}^{Y},\,\gamma\in\Gamma_{0}\setminus\bigg{(}\Gamma_{0,p}\cup\bigcup_{e_{j}\in\mathcal{E}_{\Gamma_{0}}}\Gamma_{0,e_{j}}\bigg{)}\bigg{\}}>0,$
as well as the bound
(44) $\frac{k-1}{4\pi}\sup_{\tau\in\mathcal{F}_{1}^{Y}\setminus\mathcal{F}_{1}^{k/(2\pi)}}\sum_{\gamma\in\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}=O\big{(}k^{\frac{3}{2}}\big{)}.$
By means of (42), we thus deduce from (43) and (44) in the case $Y<\frac{k}{2\pi}$ the bound
(45) $\sup_{\tau\in\mathcal{F}_{1}^{Y}}\|f(\tau)\|_{\mathrm{Pet}}^{2}\leq\sup_{\tau\in\mathcal{F}_{1}^{Y}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}=O_{\Gamma_{0},Y}\big{(}k^{\frac{3}{2}}\big{)}.$
Since $Y$ has been fixed, the claim of the theorem follows from (37), (41), and (45). ∎
Remark 3.6.
If $f\in S_{k,\chi}(\Gamma)$ is not a Hecke eigenform, then there is no evidence from the literature to suggest that the estimates (32) and (33) can be improved. Thus the estimates (32) and (33) are expected to be optimal.
4. Sup-norm bounds for Jacobi cusp forms
For $k\in\mathbb{Z}_{\geq 5}$ and $m\in\mathbb{Z}_{\geq 1}$, let $f\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$ be a Jacobi cusp form of weight $k$ and index $m$ for the full modular group $\Gamma_{0}=\mathrm{SL}_{2}(\mathbb{Z})$, which is normalized with respect to the Petersson inner product defined by (12). We now aim at bounding the quantity
$\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|f(\tau,z)\|_{\mathrm{Pet}}.$
For this we will employ the decomposition (13), namely
$f(\tau,z)=\sum_{\mu=0}^{2m-1}\varphi_{\mu}(\tau)\vartheta_{\mu,m}(\tau,z),$
where the functions $\varphi_{\mu}$ are cusp forms of weight $(k-\frac{1}{2})$ with respect to the finite index subgroup $\Gamma_{1}=\Gamma_{0}(4m)$ of $\Gamma_{0}$ and the theta functions $\vartheta_{\mu,m}$ are defined in (14).
For a fixed $\tau=\xi+i\eta\in\mathbb{H}$, consider the elliptic curve $E_{\tau}\coloneqq\mathbb{C}/\Lambda_{\tau}$ with $\Lambda_{\tau}\coloneqq\mathbb{Z}\oplus\,\tau\mathbb{Z}$. Let $O_{\tau}$ denote the identity element of $E_{\tau}$, when considered as an abelian group with $\oplus_{\tau}$ denoting the group operation, and let
$[2]\colon E_{\tau}\longrightarrow E_{\tau}$
be multiplication by $2$, given by the assignment $z\mapsto 2z\coloneqq z\oplus_{\tau}z$, which is an isogeny of degree $4$. Let $\mathcal{M}_{\tau}$ be the line bundle associated to the divisor $O_{\tau}$. Then, the theorem of the cube gives the isomorphism
$[2]^{\ast}\mathcal{M}_{\tau}\cong\mathcal{M}_{\tau}^{\otimes{4}}.$
We then put $\mathcal{L}_{\tau}\coloneqq\mathcal{M}_{\tau}^{\otimes{2}}$ and find that the theta functions $\vartheta_{\mu,m}(\tau,\cdot)$ ($\mu=0,\ldots,2m-1$) arise as global holomorphic sections of the line bundle $\mathcal{L}_{\tau}^{\otimes{m}}$.
The pointwise norm of $\vartheta_{\mu,m}(\tau,\cdot)\in H^{0}(E_{\tau},\mathcal{L}_{\tau}^{\otimes{m}})$ at the point $z=x+iy\in E_{\tau}$ (identifying $E_{\tau}$ with its universal cover $\mathbb{C}$) is given by the following formula
(46) $\|\vartheta_{\mu,m}(\tau,z)\|_{\mathcal{L}_{\tau}^{\otimes{m}}}^{2}\coloneqq|\vartheta_{\mu,m}(\tau,z)|^{2}\,\sqrt{\eta}\,e^{-\frac{4\pi my^{2}}{\eta}}.$
Let $\mu_{\mathrm{eucl}}$ denote the Euclidean metric on $E_{\tau}$; at the point $z=x+iy\in E_{\tau}$, it is given by the formula
(47) $\mu_{\mathrm{eucl}}(z)=\frac{i}{2}\cdot\frac{\mathrm{d}z\wedge\mathrm{d}\overline{z}}{\eta}=\frac{\mathrm{d}x\wedge\mathrm{d}y}{\eta}.$
Furthermore, the $\mathrm{L}^{2}$-inner product of $\vartheta_{\mu,m}(\tau,\cdot),\vartheta_{\mu^{\prime},m}(\tau,\cdot)\in H^{0}(E_{\tau},\mathcal{L}_{\tau}^{\otimes{m}})$ is given by the formula
$\langle\vartheta_{\mu,m}(\tau,\cdot),\vartheta_{\mu^{\prime},m}(\tau,\cdot)\rangle_{\mathrm{L}^{2},\mathcal{L}_{\tau}^{\otimes{m}}}$ $\coloneqq\int\limits_{E_{\tau}}\vartheta_{\mu,m}(\tau,z)\,\overline{\vartheta_{\mu^{\prime},m}(\tau,z)}\,\sqrt{\eta}\,e^{-\frac{4\pi my^{2}}{\eta}}\,\mu_{\mathrm{eucl}}(z)$
(48) $=\int\limits_{0}^{\eta}\int\limits_{0}^{1}\vartheta_{\mu,m}(\tau,z)\,\overline{\vartheta_{\mu^{\prime},m}(\tau,z)}\,\sqrt{\eta}\,e^{-\frac{4\pi my^{2}}{\eta}}\,\frac{\mathrm{d}x\wedge\mathrm{d}y}{\eta}.$
The $\mathrm{L}^{2}$-norm of $\vartheta_{\mu,m}(\tau,\cdot)\in H^{0}(E_{\tau},\mathcal{L}_{\tau}^{\otimes{m}})$ induced by the $\mathrm{L}^{2}$-inner product defined above is thus given by
(49) $\|\vartheta_{\mu,m}(\tau,\cdot)\|_{\mathrm{L}^{2},\,\mathcal{L}_{\tau}^{\otimes{m}}}^{2}=\langle\vartheta_{\mu,m}(\tau,\cdot),\vartheta_{\mu,m}(\tau,\cdot)\rangle_{\mathrm{L}^{2},\mathcal{L}_{\tau}^{\otimes{m}}}.$
We next recall the well-known fact about the orthogonality of the Jacobi theta functions (which was already implicitly used in formula (15)).
Lemma 4.1.
The set of Jacobi theta functions $\{\vartheta_{\mu,m}(\tau,\cdot)\}_{\mu=0}^{2m-1}$ constitutes an orthogonal basis for the space $H^{0}(E_{\tau},\mathcal{L}_{\tau}^{\otimes{m}})$ with respect to the $\mathrm{L}^{2}$-inner product (48).
Proof.
As above, we write $\tau=\xi+i\eta\in\mathbb{H}$ and for $\mu,\mu^{\prime}=0,\ldots,2m-1$ and $n\in\mathbb{Z}$, we define
(50) $a(\mu,n)\coloneqq n-\frac{\mu}{2m}\quad\text{and}\quad b(\mu,n)\coloneqq 2mn-\mu.$
Recalling the definition of the Jacobi theta function (14), we compute for the $\mathrm{L}^{2}$-inner product (48)
$\langle\vartheta_{\mu,m}(\tau,\cdot),\vartheta_{\mu^{\prime},m}(\tau,\cdot)\rangle_{\mathrm{L}^{2},\,\mathcal{L}_{\tau}^{\otimes{m}}}$
(51) $\qquad=\int\limits_{0}^{\eta}\sum_{n\in\mathbb{Z}}\,\sum_{n^{\prime}\in\mathbb{Z}}e^{2\pi im\tau a(\mu,n)^{2}}\overline{e^{2\pi im\tau a(\mu^{\prime},n^{\prime})^{2}}}\bigg{(}\int\limits_{0}^{1}e^{2\pi izb(\mu,n)}\overline{e^{2\pi izb(\mu^{\prime},n^{\prime})}}\,\mathrm{d}x\bigg{)}\,e^{-\frac{4\pi my^{2}}{\eta}}\,\frac{\mathrm{d}y}{\sqrt{\eta}}.$
Observing that the inequalities $-2m<\mu-\mu^{\prime}<2m$ imply that the equality
$b(\mu,n)=b(\mu^{\prime},n^{\prime})\quad\Longleftrightarrow\quad 2m(n-n^{\prime})=\mu-\mu^{\prime}$
can only hold if $n=n^{\prime}$ and $\mu=\mu^{\prime}$, we deduce
$\int\limits_{0}^{1}e^{2\pi izb(\mu,n)}\overline{e^{2\pi izb(\mu^{\prime},n^{\prime})}}\,\mathrm{d}x$ $=e^{-2\pi y(b(\mu,n)+b(\mu^{\prime},n^{\prime}))}\int\limits_{0}^{1}e^{2\pi ix(b(\mu,n)-b(\mu^{\prime},n^{\prime}))}\,\mathrm{d}x$
(52) =$=\begin{cases}e^{-4\pi yb(\mu,n)},&\text{if }\mu=\mu^{\prime}\text{ and }n=n^{\prime},\\ 0,&\text{else}.\end{cases}$
This proves the pairwise orthogonality of the $2m$ Jacobi theta functions $\{\vartheta_{\mu,m}(\tau,\cdot)\}_{\mu=0}^{2m-1}$ and thus their linear independence. Now, since we have $\deg(\mathcal{L}_{\tau}^{\otimes{m}})=2m$, the Riemann–Roch Theorem on the elliptic curve $E_{\tau}$ shows that
$\dim H^{0}(E_{\tau},\mathcal{L}_{\tau}^{\otimes{m}})=2m,$
from which the claim follows. ∎
Proposition 4.2.
With notations as above, we have for $\mu=0,\ldots,2m-1$, the bound
$\sup_{\tau\in\mathbb{H}}\|\vartheta_{\mu,m}(\tau,\cdot)\|_{\mathrm{L}^{2},\,\mathcal{L}_{\tau}^{\otimes{m}}}^{2}=O\bigg{(}\frac{1}{\sqrt{m}}\bigg{)}.$
Proof.
Let $\tau=\xi+i\eta\in\mathbb{H}$, $z=x+iy\in\mathbb{C}$, and $\mu=0,\ldots,2m-1$. For $\tau\in\mathbb{H}$, we need to bound the $\mathrm{L}^{2}$-norm (49) of $\vartheta_{\mu,m}(\tau,\cdot)\in H^{0}(E_{\tau},\mathcal{L}_{\tau}^{\otimes{m}})$, namely the quantity
$\|\vartheta_{\mu,m}(\tau,\cdot)\|_{\mathrm{L}^{2},\,\mathcal{L}_{\tau}^{\otimes{m}}}^{2}=\langle\vartheta_{\mu,m}(\tau,\cdot),\vartheta_{\mu,m}(\tau,\cdot)\rangle_{\mathrm{L}^{2},\,\mathcal{L}_{\tau}^{\otimes{m}}}.$
Recalling the notation (50) and introducing $a(n)\coloneqq a(\mu,n)$ as well as $b(n)\coloneqq b(\mu,n)$ as a short-hand, we arrive as in (51) at the equality
$\|\vartheta_{\mu,m}(\tau,\cdot)\|_{\mathrm{L}^{2},\,\mathcal{L}_{\tau}^{\otimes{m}}}^{2}$
(53) $\qquad=\int\limits_{0}^{\eta}\sum_{n\in\mathbb{Z}}\,\sum_{n^{\prime}\in\mathbb{Z}}e^{2\pi im\tau a(n)^{2}}\overline{e^{2\pi im\tau a(n^{\prime})^{2}}}\bigg{(}\int\limits_{0}^{1}e^{2\pi izb(n)}\overline{e^{2\pi izb(n^{\prime})}}\,\mathrm{d}x\bigg{)}\,e^{-\frac{4\pi my^{2}}{\eta}}\,\frac{\mathrm{d}y}{\sqrt{\eta}}.$
Since (by (52))
$\int\limits_{0}^{1}e^{2\pi izb(n)}\overline{e^{2\pi izb(n^{\prime})}}\,\mathrm{d}x=\begin{cases}e^{-4\pi yb(n)},&\text{ if }n=n^{\prime},\\ 0,&\text{ else},\end{cases}$
we arrive at
(54) $\sum_{n\in\mathbb{Z}}\,\sum_{n^{\prime}\in\mathbb{Z}}e^{2\pi im\tau a(n)^{2}}\overline{e^{2\pi im\tau a(n^{\prime})^{2}}}\int\limits_{0}^{1}e^{2\pi izb(n)}\overline{e^{2\pi izb(n^{\prime})}}\,\mathrm{d}x=\sum_{n\in\mathbb{Z}}e^{-4\pi m\eta\big{(}n-\frac{\mu}{2m}\big{)}^{2}-4\pi y(2mn-\mu)}.$
Substituting (54) into (53) and using an integral test, we find the bound
$\|\vartheta_{\mu,m}(\tau,\cdot)\|_{\mathrm{L}^{2},\,\mathcal{L}_{\tau}^{\otimes{m}}}^{2}$ $=\sum_{n\in\mathbb{Z}}\int\limits_{0}^{\eta}e^{-4\pi m\eta\big{(}n-\frac{\mu}{2m}\big{)}^{2}-4\pi y(2mn-\mu)}\,e^{-\frac{4\pi my^{2}}{\eta}}\,\frac{\mathrm{d}y}{\sqrt{\eta}}$
(55) $\leq\int\limits_{0}^{\eta}\int\limits_{-\infty}^{\infty}e^{-4\pi m\eta\big{(}\nu-\frac{\mu}{2m}\big{)}^{2}-4\pi y(2m\nu-\mu)-\frac{4\pi my^{2}}{\eta}}\,\mathrm{d}\nu\,\frac{\mathrm{d}y}{\sqrt{\eta}}$
(56) $\quad+\int\limits_{0}^{\eta}e^{-\frac{\pi\eta\mu^{2}}{m}+4\pi y\mu-\frac{4\pi my^{2}}{\eta}}\,\frac{\mathrm{d}y}{\sqrt{\eta}}.$
Now, we rewrite the exponent of the integrand in (55) in the form
$4\pi m\eta\bigg{(}\nu-\frac{\mu}{2m}\bigg{)}^{2}+4\pi y(2m\nu-\mu)+\frac{4\pi my^{2}}{\eta}$
$\qquad=\bigg{(}\sqrt{\frac{\pi\eta}{m}}(2m\nu-\mu)\bigg{)}^{2}+4\pi y(2m\nu-\mu)+\bigg{(}2\sqrt{\frac{\pi m}{\eta}}y\bigg{)}^{2}$
$\qquad=\bigg{(}\sqrt{\frac{\pi\eta}{m}}(2m\nu-\mu)+2\sqrt{\frac{\pi m}{\eta}}y\bigg{)}^{2}.$
Substituting
$\rho\coloneqq\sqrt{\frac{\pi\eta}{m}}(2m\nu-\mu)+2\sqrt{\frac{\pi m}{\eta}}y$
into (55), we obtain for the inner integral
$\int\limits_{-\infty}^{\infty}e^{-4\pi m\eta\big{(}\nu-\frac{\mu}{2m}\big{)}^{2}-4\pi y(2m\nu-\mu)-\frac{4\pi my^{2}}{\eta}}\,\mathrm{d}\nu=\frac{1}{2\sqrt{\pi m\eta}}\int\limits_{-\infty}^{\infty}e^{-\rho^{2}}\,\mathrm{d}\rho=\frac{1}{2\sqrt{m}}\frac{1}{\sqrt{\eta}}.$
From this, we compute the double integral (55) as
(57) $\int\limits_{0}^{\eta}\int\limits_{-\infty}^{\infty}e^{-4\pi m\eta\big{(}\nu-\frac{\mu}{2m}\big{)}^{2}-4\pi y(2m\nu-\mu)-\frac{4\pi my^{2}}{\eta}}\,\mathrm{d}\nu\,\frac{\mathrm{d}y}{\sqrt{\eta}}=\frac{1}{2\sqrt{m}}\frac{1}{\sqrt{\eta}}\int\limits_{0}^{\eta}\frac{\mathrm{d}y}{\sqrt{\eta}}=\frac{1}{2\sqrt{m}}.$
For the integral (56), we find in a similar way
(58) $\int\limits_{0}^{\eta}e^{-\frac{\pi\eta\mu^{2}}{m}+4\pi y\mu-\frac{4\pi my^{2}}{\eta}}\,\frac{\mathrm{d}y}{\sqrt{\eta}}\leq\int\limits_{-\infty}^{\infty}e^{-\big{(}-\sqrt{\frac{\pi\eta}{m}}\mu+2\sqrt{\frac{\pi m}{\eta}}y\big{)}^{2}}\,\frac{\mathrm{d}y}{\sqrt{\eta}}=\frac{1}{2\sqrt{m}}.$
Adding up the bounds (57) and (58), yields
$\|\vartheta_{\mu,m}(\tau,\cdot)\|_{\mathrm{L}^{2},\,\mathcal{L}_{\tau}^{\otimes{m}}}^{2}\leq\frac{1}{\sqrt{m}},$
which proves the claim. ∎
Proposition 4.3.
With notations as above, we have the bound
(59) $\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\sum_{\mu=0}^{2m-1}\|\vartheta_{\mu,m}(\tau,z)\|_{\mathcal{L}_{\tau}^{\otimes{m}}}^{2}=O(\sqrt{m}).$
Proof.
We start by considering the Bergman kernel $B_{\mathcal{L}_{\tau}^{\otimes{m}}}(z)$ associated to the line bundle $\mathcal{L}_{\tau}^{\otimes{m}}$ on the elliptic curve $E_{\tau}$ equipped with the Kähler form $\omega=\mu_{\mathrm{eucl}}$ (see formula (47)). From the bound (17), we recall
$B_{\mathcal{L}_{\tau}^{\otimes{m}}}(z)=\mathrm{det}_{\omega}\big{(}\mathrm{c}_{1}(\mathcal{L}_{\tau},\|\,\cdot\,\|_{\mathcal{L}_{\tau}})(z)\big{)}\,m+O(1),$
where $\mathrm{c}_{1}(\mathcal{L}_{\tau},\|\,\cdot\,\|_{\mathcal{L}_{\tau}})(z)$ is the curvature form of $\mathcal{L}_{\tau}$ at the point $z\in E_{\tau}$. Using equations (16) and (46), the curvature form $\mathrm{c}_{1}(\mathcal{L}_{\tau},\|\,\cdot\,\|_{\mathcal{L}_{\tau}})(z)$ is given by the formula (identifying again $E_{\tau}$ with its universal cover $\mathbb{C}$ and writing $z=x+iy$)
$\mathrm{c}_{1}(\mathcal{L}_{\tau},\|\,\cdot\,\|_{\mathcal{L}_{\tau}})(z)$ $=-\frac{i}{2\pi}\partial_{z}\partial_{\overline{z}}\log\|\vartheta_{\mu,1}(\tau,z)\|_{\mathcal{L}_{\tau}}^{2}$
$=-\frac{i}{2\pi}\partial_{z}\partial_{\overline{z}}\bigg{(}-\frac{4\pi y^{2}}{\eta}\bigg{)}=2\mu_{\mathrm{eucl}}(z)=2\omega(z).$
This yields $\mathrm{det}_{\omega}\big{(}\mathrm{c}_{1}(\mathcal{L}_{\tau},\|\,\cdot\,\|_{\mathcal{L}_{\tau}})(z)\big{)}=2$, and therefore we get
(60) $B_{\mathcal{L}_{\tau}^{\otimes{m}}}(z)=2m+O(1).$
Setting
$\Theta_{\mu,m}(\tau,z)\coloneqq\frac{\vartheta_{\mu,m}(\tau,z)}{\|\vartheta_{\mu,m}(\tau,\cdot)\|_{\mathrm{L}^{2},\mathcal{L}^{\otimes{m}}}}\qquad(\mu=0,\ldots,2m-1)$
and using Lemma 4.1, we deduce that the set $\{\Theta_{\mu,m}(\tau,\cdot)\}_{\mu=0}^{2m-1}$ constitutes an orthonormal basis for the space $H^{0}(E_{\tau},\mathcal{L}_{\tau}^{\otimes{m}})$ with respect to the $\mathrm{L}^{2}$-inner product (48). By the very definition of the Bergman kernel we thus have
$B_{\mathcal{L}_{\tau}^{\otimes{m}}}(z)=\sum_{\mu=0}^{2m-1}\|\Theta_{\mu,m}(\tau,z)\|_{\mathcal{L}_{\tau}^{\otimes{m}}}^{2}$
for any $z\in E_{\tau}$. Finally, using Proposition 4.2 and observing that
$\sum_{\mu=0}^{2m-1}\|\vartheta_{\mu,m}(\tau,z)\|_{\mathcal{L}_{\tau}^{\otimes{m}}}^{2}\leq\frac{1}{\sqrt{m}}\sum_{\mu=0}^{2m-1}\|\Theta_{\mu,m}(\tau,z)\|_{\mathcal{L}_{\tau}^{\otimes{m}}}^{2}=\frac{1}{\sqrt{m}}B_{\mathcal{L}_{\tau}^{\otimes{m}}}(z),$
the claim follows by means of (60). ∎
Theorem 4.4.
For $k\in\mathbb{Z}_{\geq 5}$ and $m\in\mathbb{Z}_{\geq 1}$, let $f\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$ be a Jacobi cusp form of weight $k$ and index $m$ for the full modular group $\Gamma_{0}=\mathrm{SL}_{2}(\mathbb{Z})$, which is normalized with respect to the Petersson inner product. Then, we have the bound
(61) $\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|f(\tau,z)\|^{2}_{\mathrm{Pet}}=O_{\epsilon}\big{(}k^{\frac{3}{2}}\,m^{2+\epsilon}\big{)},$
where the implied constant depends only on the choice of $\epsilon>0$.
Proof.
Substituting the decomposition (13) of the Jacobi form $f(\tau,z)$ into its pointwise Petersson norm and applying the Cauchy–Schwartz inequality, we compute
$\|f(\tau,z)\|^{2}_{\mathrm{Pet}}$ $=\bigg{|}\sum_{\mu=0}^{2m-1}\varphi_{\mu}(\tau)\vartheta_{\mu,m}(\tau,z)\bigg{|}^{2}\,\eta^{k}\,e^{-\frac{4\pi my^{2}}{\eta}}$
$\leq\bigg{(}\sum_{\mu=0}^{2m-1}|\varphi_{\mu}(\tau)|^{2}\,\eta^{k-1/2}\bigg{)}\bigg{(}\sum_{\mu=0}^{2m-1}|\vartheta_{\mu,m}(\tau,z)|\,\eta^{1/2}\,e^{-\frac{4\pi my^{2}}{\eta}}\bigg{)}$
$=\bigg{(}\sum_{\mu=0}^{2m-1}\|\varphi_{\mu}(\tau)\|^{2}_{\mathrm{Pet}}\bigg{)}\bigg{(}\sum_{\mu=0}^{2m-1}\|\vartheta_{\mu,m}(\tau,z)\|_{\mathcal{L}_{\tau}^{\otimes{m}}}^{2}\bigg{)}.$
Thus
(62) $\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|f(\tau,z)\|^{2}_{\mathrm{Pet}}\leq\sup_{\tau\in\mathbb{H}}\sum_{\mu=0}^{2m-1}\|\varphi_{\mu}(\tau)\|^{2}_{\mathrm{Pet}}\cdot\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\sum_{\mu=0}^{2m-1}\|\vartheta_{\mu,m}(\tau,z)\|_{\mathcal{L}_{\tau}^{\otimes{m}}}^{2}.$
Since $f$ is normalized with respect to the Petersson inner product, we have from (15) that
(63) $\langle f,f\rangle_{\mathrm{Pet}}=\frac{1}{\sqrt{4m}\,[\Gamma_{0}:\Gamma_{1}]}\sum_{\mu=0}^{2m-1}\langle\varphi_{\mu},\varphi_{\mu}\rangle_{\mathrm{Pet}}=1.$
Combining (63) with the bound (35) given in Theorem 3.5, we find the bound
(64) $\sum_{\mu=0}^{2m-1}\|\varphi_{\mu}(\tau)\|^{2}_{\mathrm{Pet}}=\sum_{\mu=0}^{2m-1}\langle\varphi_{\mu},\varphi_{\mu}\rangle_{\mathrm{Pet}}\,\frac{\|\varphi_{\mu}(\tau)\|^{2}_{\mathrm{Pet}}}{\langle\varphi_{\mu},\varphi_{\mu}\rangle_{\mathrm{Pet}}}\leq C\,\sqrt{4m}\,[\Gamma_{0}:\Gamma_{1}]\,k^{\frac{3}{2}},$
where $C$ is a positive constant depending on $\Gamma_{0}$. Recalling now that
(65) $[\Gamma_{0}:\Gamma_{1}]=[\mathrm{SL}_{2}(\mathbb{Z}):\Gamma_{0}(4m)]=O_{\epsilon}\big{(}m^{1+\epsilon}\big{)},$
where the implied constant depends only on the choice of $\epsilon>0$, we derive from the bounds (64) and (65), the bound
(66) $\sup_{\tau\in\mathbb{H}}\sum_{\mu=0}^{2m-1}\|\varphi_{\mu}(\tau)\|^{2}_{\mathrm{Pet}}=O_{\epsilon}\big{(}k^{\frac{3}{2}}\,m^{\frac{3}{2}+\epsilon}\big{)},$
where the implied constant depends only on the choice of $\epsilon>0$. The claimed bound (61) finally follows from (62) by combining the bound (66) with the bound (59) established in Proposition 4.3. ∎
Remark 4.5.
The bound (61) is polynomial in $k$ and $m$, and thus improves W. Kohnen’s bound (1), which is exponential in $k$. Comparing our bound with the one obtained by P. Anamby and S. Das in [AD23], there is an improvement with regard to the polynomial growth in $k$, while the polynomial growth in $m$ is slightly worse.
Acknowledgements
The first-named author expresses his gratitude to the Departments of Mathematics of Technical University of Darmstadt and of Humboldt University of Berlin for their hospitality, where this article was realized. The first-named author also acknowledges support from the INSPIRE research grant DST/INSPIRE/04/2015/002263 and the SERB Matrics grant MTR/2018/000636. The first-named author also would like to thank Prof. D. S. Nagaraj for helpful discussions. The second-named author acknowledges support from the DFG Cluster of Excellence MATH+. The third-named author acknowledges support from the LOEWE research unit Uniformized Structures in Arithmetic and Geometry. Lastly, all the three authors would like to acknowledge the hospitality of the International Centre for Theoretical Sciences (ICTS) during a visit for participating in the program Algebraic and Analytic Aspects of Automorphic Forms (Code: ICTS/aforms2019/02), where this project was initiated.
Finally, the authors express their gratitude to Edgar Assing, whose observations have helped to improve our initial bound (3) by a factor of $\sqrt{m}$.

### v3

@@
-Sup-norm bounds for Jacobi cusp forms
+
 Abstract.
-In this article, we give bounds for the natural invariant norm of cusp forms of real weight $k$ and character $\chi$ for any cofinite Fuchsian subgroup $\Gamma\subset\mathrm{SL}_{2}(\mathbb{R})$.
-Using the representation of Jacobi cusp forms of integral weight $k$ and index $m$ for the modular group $\Gamma_{0}=\mathrm{SL}_{2}(\mathbb{Z})$ as linear combinations of modular forms of weight $k-\frac{1}{2}$ for some congruence subgroup of $\Gamma_{0}$ (depending on $m$) and suitable Jacobi theta functions, we derive bounds for the natural invariant norm of these Jacobi cusp forms.
-More specifically, letting $J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$ denote the complex vector space of Jacobi cusp forms under consideration and $\|\cdot\|_{\mathrm{Pet}}$ the pointwise Petersson norm on $J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$, we prove that for given $\epsilon>0$, the bound
-$\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|f(\tau,z)\|_{\mathrm{Pet}}=O_{\epsilon}\big{(}k^{\frac{3}{4}}m^{1+\epsilon}\big{)}$
-holds for any $f\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$, which is normalized with respect to the Petersson inner product, where the implied constant depends only on the choice of $\epsilon>0$.
+In this article, we give $L^{\infty}$-norm bounds for the natural invariant norm of cusp forms of real weight $k$ and character $\chi$ for any cofinite Fuchsian subgroup $\Gamma\subset\mathrm{SL}_{2}(\mathbb{R})$.
+Using the representation of Jacobi cusp forms of integral weight $k$ and index $m$ for the modular group $\Gamma_{0}=\mathrm{SL}_{2}(\mathbb{Z})$ as linear combinations of modular forms of weight $k-\frac{1}{2}$ for some congruence subgroup of $\Gamma_{0}$ (depending on $m$) and suitable Jacobi theta functions, we derive $L^{\infty}$-norm bounds for the natural invariant norm of these Jacobi cusp forms.
+More specifically, letting $J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$ denote the complex vector space of Jacobi cusp forms under consideration and $\|\cdot\|_{\mathrm{Pet}}$ the pointwise Petersson norm on $J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$, we prove that for $k\in\mathbb{Z}_{\geq 5}$ and $m\in\mathbb{Z}_{\geq 1}$, and a given $\epsilon>0$, the $L^{\infty}$-norm bound
+$\|\phi\|_{L^{\infty}}=\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|\phi(\tau,z)\|_{\mathrm{Pet}}=O_{\Gamma_{0},\epsilon}\big{(}k\,m^{\frac{7}{4}+\epsilon}\big{)}$
+holds for any $\phi\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$, which is $L^{2}$-normalized with respect to the Petersson inner product, where the implied constant depends on $\Gamma_{0}$ and the choice of $\epsilon>0$.
 2010 Mathematics Subject Classification:
@@
 Jorgenson, and J.
-Kramer obtained optimal sup-norm bounds on average for cusp forms of even weight $k$ for any cofinite Fuchsian subgroup $\Gamma\subset\mathrm{SL}_{2}(\mathbb{R})$.
-These bounds turn out to be uniform with respect to the subgroup $\Gamma$.
-Moreover, in [FJK19], effective versions for these sup-norm bounds are given.
-With regard to sup-norm bounds for individual Hecke eigenforms of large level, we mention, for example, the results by V.
+Kramer obtained optimal $L^{\infty}$-norm bounds on average for cusp forms of even weight $k$ for any cofinite Fuchsian subgroup $\Gamma\subset\mathrm{SL}_{2}(\mathbb{R})$.
+These $L^{\infty}$-norm bounds turn out to be uniform with respect to the subgroup $\Gamma$.
+Moreover, in [FJK19], effective versions for these $L^{\infty}$-norm bounds are given.
+With regard to $L^{\infty}$-norm bounds for individual Hecke eigenforms of large level, we mention, for example, the results by V.
 Blomer and R.
 Holowinsky in [BH11].
-So far, less attention has been devoted to the study of sup-norm bounds for Jacobi cusp forms.
+So far, less attention has been devoted to the study of $L^{\infty}$-norm bounds for Jacobi cusp forms.
 The first comprehensive study of Jacobi forms was undertaken by M.
@@
 We denote the complex vector space of Jacobi cusp forms of weight $k$ and index $m$ for $\Gamma_{0}$ by $J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$.
-The pointwise Petersson norm of a Jacobi form $f\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$ is then defined by
-$\|f(\tau,z)\|^{2}_{\mathrm{Pet}}\coloneqq|f(\tau,z)|^{2}\,\mathrm{Im}(\tau)^{k}\,e^{-\frac{4\pi m\mathrm{Im}(z)^{2}}{\mathrm{Im}(\tau)}}\qquad(\tau\in\mathbb{H},\,z\in\mathbb{C}).$
-Let $F$ be a Siegel cusp form of weight $k$ for the Siegel modular group $\mathrm{Sp}_{4}(\mathbb{Z})$, and let $\{f_{m}\}_{m\geq 1}$ be the set of Jacobi forms appearing in the Fourier–Jacobi expansion of $F$, i.
-e., $f_{m}\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$.
+The pointwise Petersson norm of a Jacobi form $\phi\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$ is then defined by
+$\|\phi(\tau,z)\|^{2}_{\mathrm{Pet}}\coloneqq|\phi(\tau,z)|^{2}\,\mathrm{Im}(\tau)^{k}\,e^{-\frac{4\pi m\mathrm{Im}(z)^{2}}{\mathrm{Im}(\tau)}}\qquad(\tau\in\mathbb{H},\,z\in\mathbb{C}).$
+Let $F$ be a Siegel cusp form of weight $k$ for the Siegel modular group $\mathrm{Sp}_{4}(\mathbb{Z})$, and let $\{\phi_{m}\}_{m\geq 1}$ be the set of Jacobi cusp forms appearing in the Fourier–Jacobi expansion of $F$, i.
+e., $\phi_{m}\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$.
 Then, for any $\epsilon>0$, W.
-Kohnen proved the following sup-norm bound for the pointwise Petersson norm of $f_{m}$ in [Ko93]
-(1) $\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|f_{m}(\tau,z)\|_{\mathrm{Pet}}=O_{F,\epsilon}\big{(}m^{\frac{k}{2}-\frac{2}{9}+\epsilon}\big{)},$
-where the implied constant depends on the Siegel cusp form $F$ and on the choice of $\epsilon>0$. Motivated by the Ramanujan–Petersson conjecture, W. Kohnen then conjectured the bound
-$\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|f_{m}(\tau,z)\|_{\mathrm{Pet}}=O_{F,\epsilon}\big{(}m^{\frac{k-1}{2}+\epsilon}\big{)},$
-where the implied constant depends on the Siegel cusp form $F$ and on the choice of $\epsilon>0$.
+Kohnen proved the following $L^{\infty}$-norm bound for the pointwise Petersson norm of $\phi_{m}$ in [Ko93]
+(1) $\|\phi_{m}\|_{L^{\infty}}=\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|\phi_{m}(\tau,z)\|_{\mathrm{Pet}}=O_{F,\epsilon}\big{(}m^{\frac{k}{2}-\frac{2}{9}+\epsilon}\big{)},$
+where the implied constant depends on the Siegel cusp form $F$ and the choice of $\epsilon>0$. Motivated by the Ramanujan–Petersson conjecture, W. Kohnen then conjectured the $L^{\infty}$-norm bound
+$\|\phi_{m}\|_{L^{\infty}}=\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|\phi_{m}(\tau,z)\|_{\mathrm{Pet}}=O_{F,\epsilon}\big{(}m^{\frac{k-1}{2}+\epsilon}\big{)},$
+where the implied constant depends on the Siegel cusp form $F$ and the choice of $\epsilon>0$.
 More recently, P.
 Anamby and S.
-Das established in [AD23] a general sup-norm bound for the pointwise Petersson norm of $f\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$, which is normalized with respect to the Petersson inner product, i.
+Das established in [AD23] a general $L^{\infty}$-norm bound for the pointwise Petersson norm of any Jacobi cusp form $\phi\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$, which is $L^{2}$-normalized with respect to the Petersson inner product, i.
 e., for which we have
-$\int\limits\limits_{\Gamma_{0}\ltimes\mathbb{Z}^{2}\backslash\mathbb{H}\times\mathbb{C}}\|f(\tau,z)\|^{2}_{\mathrm{Pet}}\,\frac{\mathrm{d}\xi\wedge\mathrm{d}\eta\wedge\mathrm{d}x\wedge\mathrm{d}y}{\eta^{3}}=1\qquad(\tau=\xi+i\eta,\,z=x+iy).$
-Their bound is (see Theorem 1.4 in[AD23])
-(2) $\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|f(\tau,z)\|_{\mathrm{Pet}}=O_{\epsilon}\big{(}k\,m\big{)},$
+$\|\phi\|^{2}_{L^{2}}=\int\limits\limits_{\Gamma_{0}\ltimes\mathbb{Z}^{2}\backslash\mathbb{H}\times\mathbb{C}}\|\phi(\tau,z)\|^{2}_{\mathrm{Pet}}\,\frac{\mathrm{d}\xi\wedge\mathrm{d}\eta\wedge\mathrm{d}x\wedge\mathrm{d}y}{\eta^{3}}=1\qquad(\tau=\xi+i\eta,\,z=x+iy).$
+Their $L^{\infty}$-norm bound is (see Theorem 1.4 in[AD23])
+(2) $\|\phi\|_{L^{\infty}}=\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|\phi(\tau,z)\|_{\mathrm{Pet}}=O_{\epsilon}\big{(}(k\,m)^{1+\epsilon}\big{)},$
 where the implied constant depends on the choice of $\epsilon>0$.
 1.2. Main results
-The goal of this article is to provide new sup-norm bounds for the pointwise Petersson norm for Jacobi forms of integral weight $k$ and integral index $m$ for $\Gamma_{0}$, which are normalized with respect to the Petersson inner product.
-The main result in this respect is given in Theorem 4.4 and states for $f\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$, which is normalized with respect to the Petersson inner product, that
-(3) $\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|f(\tau,z)\|_{\mathrm{Pet}}=O_{\epsilon}\big{(}k^{\frac{3}{4}}\,m^{1+\epsilon}\big{)},$
-where the implied constant depends only on the choice of $\epsilon>0$.
-For the proof, we essentially use the representation of the Jacobi cusp forms under consideration as linear combinations of modular forms of weight $k-\frac{1}{2}$ for some congruence subgroup of $\Gamma_{0}$ (depending on $m$) and suitable Jacobi theta functions; we then need to derive bounds for the pointwise Petersson norms of these functions to arrive at our result.
-Comparing our bound (3) with the bound (2) by P.
+The goal of this article is to provide $L^{\infty}$-norm bounds for the pointwise Petersson norm for Jacobi forms of weight $k\in\mathbb{Z}_{\geq 5}$ and index $m\in\mathbb{Z}_{\geq 1}$ for $\Gamma_{0}$, which are $L^{2}$-normalized.
+The main result in this respect is given in Theorem 4.4 and states for any $L^{2}$-normalized Jacobi cusp form $\phi\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$ the $L^{\infty}$-norm bound
+(3) $\|\phi\|_{L^{\infty}}=\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|\phi(\tau,z)\|_{\mathrm{Pet}}=O_{\Gamma_{0},\epsilon}\big{(}k\,m^{\frac{7}{4}+\epsilon}\big{)}$
+holds, where the implied constant depends on $\Gamma_{0}$ and the choice of $\epsilon>0$.
+For the proof, we essentially use the representation of the Jacobi cusp forms under consideration as linear combinations of modular forms of weight $k-\frac{1}{2}$ for some congruence subgroup of $\Gamma_{0}$ (depending on $m$) and suitable Jacobi theta functions; we then need to derive bounds for the $L^{\infty}$-norms of these functions to arrive at our result.
+Comparing our $L^{\infty}$-norm bound (3) with the $L^{\infty}$-norm bound (2) by P.
 Anamby and S.
-Das, we realize an improvement with regard to the polynomial growth in $k$ by a factor of $k^{\frac{1}{4}}$, while the polynomial growth of the bound (2) in $m$ is better by a factor of $m^{\epsilon}$ (for any $\epsilon>0$).
-In order to be able to derive our bound (3), we need sup-norm bounds for the pointwise Petersson norm of cusp forms of positive real weight $k$ and character $\chi$ for any cofinite Fuchsian subgroup $\Gamma$ of $\mathrm{SL}_{2}(\mathbb{R})$.
-Such bounds have been derived for finite index subgroups of $\mathrm{SL}_{2}(\mathbb{Z})$ by Steiner in [St16].
-Although these bounds are uniform with respect to the index of the finite index subgroups of $\mathrm{SL}_{2}(\mathbb{Z})$, there are mild restrictions for a direct application of the bounds derived in [St16] to our setting.
-On the other hand, such bounds could also be derived from [FJK16] with some extra work.
+Das, shows that the polynomial growth in $k$ is the same, while the polynomial growth of their bound in $m$ is better by a factor of $m^{\frac{3}{4}}$.
+In order to be able to derive our $L^{\infty}$-norm bound (3), we need $L^{\infty}$-norm bounds for the pointwise Petersson norm of cusp forms of positive real weight $k$ and character $\chi$ for any cofinite Fuchsian subgroup $\Gamma$ of $\mathrm{SL}_{2}(\mathbb{R})$.
+Such $L^{\infty}$-norm bounds have been derived for finite index subgroups of $\mathrm{SL}_{2}(\mathbb{Z})$ by Steiner in [St16].
+Although these $L^{\infty}$-norm bounds are uniform with respect to the index of the finite index subgroups of $\mathrm{SL}_{2}(\mathbb{Z})$, there are mild restrictions for a direct application of the $L^{\infty}$-norm bounds derived in [St16] to our setting.
+On the other hand, such $L^{\infty}$-norm bounds could also be derived from [FJK16] with some extra work.
 However, we provide here new, alternative proofs for the results of [FJK16] applying to any positive real weight $k$ and any character $\chi$ by using the Bergman kernel for the modular curve associated to $\Gamma$.
@@
 (4) $\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}=\sum\limits_{j=1}^{d_{k}}\|f_{j}(\tau)\|^{2}_{\mathrm{Pet}}.$
-As a second main result of this article, we establish in Theorem 3.3, assuming that $k\in\mathbb{R}_{\geq 5}$, for $\Gamma$ being cocompact without elliptic elements the bound
-$\sup_{z\in\mathbb{H}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}=O_{\Gamma}(k);$
-moreover, for $\Gamma$ being cofinite, we give the bound
-$\sup_{z\in\mathbb{H}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}=O_{\Gamma}\big{(}k^{\frac{3}{2}}\big{)},$
+As a second main result of this article, we establish in Theorem 3.3, assuming that $k\in\mathbb{R}_{\geq 5}$, for $\Gamma$ being cocompact without elliptic elements the $L^{\infty}$-norm bound
+$\|B_{k,\chi}\|_{L^{\infty}}=\sup_{\tau\in\mathbb{H}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}=O_{\Gamma}(k);$
+moreover, for $\Gamma$ being cofinite, we give the $L^{\infty}$-norm bound
+$\|B_{k,\chi}\|_{L^{\infty}}=\sup_{z\in\mathbb{H}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}=O_{\Gamma}\big{(}k^{\frac{3}{2}}\big{)},$
 where the implied constants depend only on the Fuchsian subgroup $\Gamma$.
-Due to the relation (4), these results reprove the sup-norm bounds on average obtained in [FJK16], but now for any real weight $k\in\mathbb{R}_{\geq 5}$ and any character $\chi$.
-Based on these results, we are then also able to prove the uniformity of the above bounds with respect to the subgroup $\Gamma$ in Theorem 3.5.
+Due to the relation (4), these results reprove the ${L^{\infty}}$-norm bounds on average obtained in [FJK16], but now for any real weight $k\in\mathbb{R}_{\geq 5}$ and any character $\chi$.
+Based on these results, we are then also able to prove the uniformity of the above ${L^{\infty}}$-norm bounds with respect to the subgroup $\Gamma$ in Theorem 3.5.
 1.3. Outline
@@
 In the subsequent, second section we collect all the necessary prerequisites for the sequel of the paper.
-In particular, we introduce the definitions of cusp forms and Jacobi cusp forms together with their respective (pointwise) Petersson inner products.
+In particular, we introduce the definitions of cusp forms and Jacobi cusp forms together with their respective pointwise Petersson norms.
 Furthermore, we define the Bergman kernel for modular curves and state its basic properties.
-We close the section by recalling asymptotics of Bergman kernels associated to tensor powers of holomorphic line bundles on compact complex Kähler manifolds due to [Bo96], which are crucial in the derivation of the bound (3).
-The third section is devoted to the revisiting of the sup-norm bounds on average obtained in [FJK16], but now for any real weight $k\in\mathbb{R}_{\geq 5}$ and any character $\chi$.
+The third section is devoted to the revisiting of the ${L^{\infty}}$-norm bounds on average obtained in [FJK16], but now for any real weight $k\in\mathbb{R}_{\geq 5}$ and any character $\chi$.
 Here, the proofs of Theorem 3.3 and Theorem 3.5 are provided.
-In the fourth section, the bound (3) is proven in Theorem 4.4.
-In addition to some straightforward bounds established for the relative $\mathrm{L}^{2}$-norm of classical Jacobi theta functions, the above mentioned asymptotics of Bergman kernels associated to tensor powers of holomorphic line bundles corresponding to these Jacobi theta functions are crucial in the derivation of the proof of Theorem 4.4.
+In the fourth section, the ${L^{\infty}}$-norm bound (3) is proven in Theorem 4.4.
+In order to establish this result, we combine the results of the third section applied to modular forms of half-integral weight with $L^{\infty}$-norm bounds of classical Jacobi theta functions, which are derived by directly estimating their defining series expansions.
 2. Preliminaries
@@
 (7) $\mathcal{F}_{Y}\coloneqq\mathrm{cl}\big{(}\mathcal{F}_{\Gamma}\setminus\big{(}\mathcal{F}^{Y}_{1}\cup\ldots\cup\mathcal{F}^{Y}_{s}\big{)}\big{)},$
-which is compact; we note that $\mathcal{F}_{Y}=\mathcal{F}_{\Gamma}$, if $\Gamma$ is cocompact. We choose $0<m_{Y}<M_{Y}$ such that for all $\tau\in\mathcal{F}_{Y}$ the inequalities
+which is compact.
+For $Y$ being sufficiently large, we note that $\mathcal{F}_{Y}=\mathcal{F}_{\Gamma}$, if $\Gamma$ is cocompact.
+We choose $0<m_{Y}<M_{Y}$ such that for all $\tau\in\mathcal{F}_{Y}$ the inequalities
 $m_{Y}\leq\mathrm{Im}(\sigma_{\mathcal{P},j}^{-1}\tau)\leq M_{Y}$
@@
 Finally, the injectivity radius $r_{\Gamma}$ is defined by
-(8) $r_{\Gamma}\coloneqq\inf\bigg{\{}\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)\,\bigg{|}\,\tau\in\mathcal{F}_{\Gamma},\,\gamma\in\Gamma\setminus\bigg{(}\bigcup_{j=1}^{s}\Gamma_{p_{j}}\cup\bigcup_{j=1}^{t}\Gamma_{e_{j}}\bigg{)}\bigg{\}}.$
+(8) $r_{\Gamma}\coloneqq\inf\bigg{\{}\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)\,\bigg{|}\,\tau\in\mathcal{F}_{\Gamma},\,\gamma\in\Gamma\setminus\bigg{(}\bigcup_{j=1}^{s}\Gamma_{p_{j}}\cup\bigcup_{j=1}^{t}\Gamma_{e_{j}}\cup\{\mathrm{id}\}\bigg{)}\bigg{\}}.$
 We note that if $X_{\Gamma}$ is compact without elliptic fixed points, i.
@@
 where
-$\mathcal{N}_{\Gamma}(\tau;\rho)\coloneqq\bigg{\{}\gamma\in\Gamma\setminus\bigg{(}\bigcup_{j=1}^{s}\Gamma_{p_{j}}\cup\bigcup_{j=1}^{t}\Gamma_{e_{j}}\bigg{)}\,\bigg{|}\,\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)\leq\rho\bigg{\}}.$
+$\mathcal{N}_{\Gamma}(\tau;\rho)\coloneqq\bigg{\{}\gamma\in\Gamma\setminus\bigg{(}\bigcup_{j=1}^{s}\Gamma_{p_{j}}\cup\bigcup_{j=1}^{t}\Gamma_{e_{j}}\cup\{\mathrm{id}\}\bigg{)}\,\bigg{|}\,\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)\leq\rho\bigg{\}}.$
 Let now $f$ be a positive, smooth, and decreasing function on $\mathbb{R}_{\geq 0}$.
@@
 For $k,m\in\mathbb{N}$, we let $J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$ denote the space of Jacobi cusp forms of weight $k$ and index $m$ for $\Gamma_{0}=\mathrm{SL}_{2}(\mathbb{Z})$, i.
-e., the space of holomorphic functions $f\colon\mathbb{H}\times\mathbb{C}\rightarrow\mathbb{C}$, which have the transformation behaviour
-$f\Big{(}\frac{a\tau+b}{c\tau+d},\frac{z+\lambda\tau+\mu}{c\tau+d}\Big{)}(c\tau+d)^{-k}\exp\Big{(}2\pi im\Big{(}\lambda^{2}\tau+2\lambda z-\frac{c(z+\lambda\tau+\mu)^{2}}{c\tau+d}\Big{)}\Big{)}=f(\tau,z)$
+e., the space of holomorphic functions $\phi\colon\mathbb{H}\times\mathbb{C}\rightarrow\mathbb{C}$, which have the transformation behaviour
+$\phi\Big{(}\frac{a\tau+b}{c\tau+d},\frac{z+\lambda\tau+\mu}{c\tau+d}\Big{)}(c\tau+d)^{-k}\exp\Big{(}2\pi im\Big{(}\lambda^{2}\tau+2\lambda z-\frac{c(z+\lambda\tau+\mu)^{2}}{c\tau+d}\Big{)}\Big{)}=\phi(\tau,z)$
 for all $\big{[}\big{(}\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\big{)},(\lambda,\mu)\big{]}\in\Gamma_{0}\ltimes\mathbb{Z}^{2}$, and which have a Fourier expansion of the form
-$f(\tau,z)=\sum\limits_{\begin{subarray}{c}n\in\mathbb{N},\,r\in\mathbb{Z}\\ 4mn-r^{2}>0\end{subarray}}c(n,r)\,q^{n}\zeta^{r}\qquad(q=e^{2\pi i\tau},\,\zeta=e^{2\pi iz}).$
-Given $f\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$, we define
-$\|f(\tau,z)\|_{\mathrm{Pet}}^{2}\coloneqq|f(\tau,z)|^{2}\,\eta^{k}\,e^{-\frac{4\pi my^{2}}{\eta}}\qquad(\tau=\xi+i\eta,\,z=x+iy),$
-which defines a $\Gamma_{0}\ltimes\mathbb{Z}^{2}$-invariant function on $\mathbb{H}\times\mathbb{C}$ called the pointwise Petersson norm of $f$.
+$\phi(\tau,z)=\sum\limits_{\begin{subarray}{c}n\in\mathbb{N},\,r\in\mathbb{Z}\\ 4mn-r^{2}>0\end{subarray}}c(n,r)\,q^{n}\zeta^{r}\qquad(q=e^{2\pi i\tau},\,\zeta=e^{2\pi iz}).$
+Given $\phi\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$, we define
+$\|\phi(\tau,z)\|_{\mathrm{Pet}}^{2}\coloneqq|\phi(\tau,z)|^{2}\,\eta^{k}\,e^{-\frac{4\pi my^{2}}{\eta}}\qquad(\tau=\xi+i\eta,\,z=x+iy),$
+which defines a $\Gamma_{0}\ltimes\mathbb{Z}^{2}$-invariant function on $\mathbb{H}\times\mathbb{C}$ called the pointwise Petersson norm of $\phi$.
 Let $\mathcal{D}_{\Gamma_{0}}$ denote a fundamental domain of the quotient space $Y_{\Gamma_{0}}\coloneqq\Gamma_{0}\ltimes\mathbb{Z}^{2}\backslash\mathbb{H}\times\mathbb{C}$, which is a $2$-dimensional complex manifold.
 The space $J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$ is equipped with the Petersson inner product
-(12) $\langle f_{1},f_{2}\rangle_{\mathrm{Pet}}\coloneqq\int\limits\limits_{\mathcal{D}_{\Gamma_{0}}}f_{1}(\tau,z)\overline{f_{2}(\tau,z)}\,\eta^{k}\,e^{-\frac{4\pi my^{2}}{\eta}}\,\frac{\mathrm{d}\xi\wedge\mathrm{d}\eta\wedge\mathrm{d}x\wedge\mathrm{d}y}{\eta^{3}}\qquad(f_{1},f_{2}\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})).$
-For $f\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$, one has the decomposition
-(13) $f(\tau,z)=\sum_{\mu=0}^{2m-1}\varphi_{\mu}(\tau)\vartheta_{\mu,m}(\tau,z),$
-where the function $\varphi_{\mu}$ is a cusp form of weight $(k-\frac{1}{2})$ for the finite index subgroup $\Gamma_{1}\coloneqq\Gamma_{0}(4m)$ of $\Gamma_{0}$, and $\vartheta_{\mu,m}$ is the Jacobi theta function
+(12) $\langle\phi_{1},\phi_{2}\rangle_{\mathrm{Pet}}\coloneqq\int\limits\limits_{\mathcal{D}_{\Gamma_{0}}}\phi_{1}(\tau,z)\overline{\phi_{2}(\tau,z)}\,\eta^{k}\,e^{-\frac{4\pi my^{2}}{\eta}}\,\frac{\mathrm{d}\xi\wedge\mathrm{d}\eta\wedge\mathrm{d}x\wedge\mathrm{d}y}{\eta^{3}}\qquad(\phi_{1},\phi_{2}\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})).$
+For $\phi\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$, one has the decomposition
+(13) $\phi(\tau,z)=\sum_{\mu=0}^{2m-1}h_{\mu}(\tau)\vartheta_{\mu,m}(\tau,z),$
+where the function $h_{\mu}$ by Remark 2 on p. 287 of [Kr86] is a cusp form of weight $(k-\frac{1}{2})$ for the finite index subgroup
+$\Gamma_{1}\coloneqq\Gamma_{0,1}(4m)=\bigg{\{}\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\Gamma_{0}\,\bigg{|}\,c\equiv 0,\,d\equiv 1\mod 4m\bigg{\}}\subseteq\Gamma_{0}(4m),$
+and $\vartheta_{\mu,m}$ is the Jacobi theta function
 (14) $\vartheta_{\mu,m}(\tau,z)\coloneqq\sum_{n\in\mathbb{Z}}e^{2\pi im\tau\big{(}n-\frac{\mu}{2m}\big{)}^{2}+2\pi iz(2mn-\mu)}.$
-As we will see below, the theta functions $\vartheta_{\mu,m}$ ($\mu=0,\ldots,2m-1$) arise for fixed $\tau\in\mathbb{H}$ as global sections of a suitable line bundle on the elliptic curve associated to $\tau$.
 In fact, it is shown in Theorem 5.1 of [EZ85] that the decomposition (14) gives rise to the isomorphism
@@
 Let now
-$f_{1}(\tau,z)=\sum_{\mu=0}^{2m-1}\varphi_{\mu,1}(\tau)\vartheta_{\mu,m}(\tau,z)\quad\text{and}\quad f_{2}(\tau,z)=\sum_{\mu=0}^{2m-1}\varphi_{\mu,2}(\tau)\vartheta_{\mu,m}(\tau,z)$
+$\phi_{1}(\tau,z)=\sum_{\mu=0}^{2m-1}h_{\mu,1}(\tau)\vartheta_{\mu,m}(\tau,z)\quad\text{and}\quad\phi_{2}(\tau,z)=\sum_{\mu=0}^{2m-1}h_{\mu,2}(\tau)\vartheta_{\mu,m}(\tau,z)$
 be two Jacobi cusp forms of weight $k$ and index $m$ for $\Gamma_{0}$. Then, the decomposition (13) gives rise to the equality (see Theorem 5.3 in [EZ85])
-(15) $\langle f_{1},f_{2}\rangle_{\mathrm{Pet}}=\frac{1}{\sqrt{4m}}\int\limits_{\mathcal{F}_{\Gamma_{0}}}\sum_{\mu=0}^{2m-1}\varphi_{\mu,1}(\tau)\overline{\varphi_{\mu,2}(\tau)}\,\eta^{k-\frac{1}{2}}\,\frac{\mathrm{d}\xi\wedge\mathrm{d}\eta}{\eta^{2}},$
-where we recall that $\mathcal{F}_{\Gamma_{0}}$ is a fundamental domain for the quotient space $X_{\Gamma_{0}}=\Gamma_{0}\backslash\mathbb{H}$.
-2.6. Asymptotics for Bergman kernels
-In this subsection, we recall asymptotics of Bergman kernels associated to tensor powers of holomorphic line bundles on compact complex Kähler manifolds, which are used in section 4 to derive bounds for theta functions.
-Let $(M,\omega)$ be a compact complex Kähler manifold of dimension $n$ with positive closed $(1,1)$-form $\omega$.
-Let $\mathcal{L}$ be a positive hermitian holomorphic line bundle on $M$ and let $H^{0}(M,\mathcal{L}^{\otimes{m}})$ denote the vector space of global holomorphic sections its $m$-th tensor power $\mathcal{L}^{\otimes{m}}$ for $m\in\mathbb{Z}_{\geq 1}$.
-Let $|\,\cdot\,|_{\mathcal{L}^{\otimes{m}}}$ and $\langle\,\cdot\,,\,\cdot\,\rangle_{\mathrm{L}^{2},\,\mathcal{L}^{\otimes{m}}}$ denote the pointwise hermitian metric and the $\mathrm{L}^{2}$-inner product on $H^{0}(M,\mathcal{L}^{\otimes{m}})$, respectively.
-Let $\{s_{j}\}$ denote an orthonormal basis of $H^{0}(M,\mathcal{L}^{\otimes{m}})$ with respect to the $\mathrm{L}^{2}$-inner product. For any $z\in M$, the function
-$B_{\mathcal{L}^{\otimes{m}}}(z)\coloneqq\sum_{j}|s_{j}(z)|_{\mathcal{L}^{\otimes{m}}}^{2}$
-is called the Bergman kernel associated to the line bundle $\mathcal{L}^{\otimes{m}}$.
-We note that the above definition is independent on the choice of an orthonormal basis of $H^{0}(M,\mathcal{L}^{\otimes{m}})$.
-Let
-(16) $\mathrm{c}_{1}(\mathcal{L},|\,\cdot\,|_{\mathcal{L}})(z)\coloneqq-\frac{i}{2\pi}\partial_{z}\partial_{\overline{z}}\log|s(z)|_{\mathcal{L}}^{2}$
-denote the curvature form of the line bundle $\mathcal{L}$ at the point $z\in M$, where $s$ is any meromorphic section of $\mathcal{L}$.
-At any $z\in M$, there exists a coordinate chart around the point $z$ such that
-$\omega(z)=\sum_{j=1}^{n}\frac{i}{2}\cdot\mathrm{d}z_{j}\wedge\mathrm{d}\overline{z}_{j}\quad\text{and}\quad\mathrm{c}_{1}(\mathcal{L},|\,\cdot\,|_{\mathcal{L}})(z)=\sum_{j=1}^{n}\frac{i}{2}\cdot\alpha_{j}\cdot\mathrm{d}z_{j}\wedge\mathrm{d}\overline{z}_{j}.$
-The complex numbers $\alpha_{1},\ldots,\alpha_{n}$ are called the eigenvalues of the curvature form $\mathrm{c}_{1}(\mathcal{L},|\,\cdot\,|_{\mathcal{L}})(z)$ at the point $z\in M$. We set
-$\mathrm{det}_{\omega}\big{(}\mathrm{c}_{1}(\mathcal{L},|\,\cdot\,|_{\mathcal{L}})(z)\big{)}\coloneqq\prod_{j=1}^{n}\alpha_{j}.$
-Since the line bundle $\mathcal{L}$ is positive, we have $\alpha_{j}>0$ for $j=1,\ldots,n$. Finally, we recall from Theorem 2.1 in [Bo96] the bound
-(17) $B_{\mathcal{L}^{\otimes{m}}}(z)=\mathrm{det}_{\omega}\big{(}\mathrm{c}_{1}(\mathcal{L},|\,\cdot\,|_{\mathcal{L}})(z)\big{)}m^{n}+O(m^{n-1}),$
-provided that $\mathcal{L}$ is a positive line bundle for any $z\in M$.
-3. Sup-norm bounds for cusp forms revisited
+(15) $\langle\phi_{1},\phi_{2}\rangle_{\mathrm{Pet}}=\frac{1}{\sqrt{4m}}\int\limits_{\mathcal{F}_{\Gamma_{0}}}\sum_{\mu=0}^{2m-1}h_{\mu,1}(\tau)\overline{h_{\mu,2}(\tau)}\,\eta^{k-\frac{1}{2}}\,\frac{\mathrm{d}\xi\wedge\mathrm{d}\eta}{\eta^{2}},$
+where $\mathcal{F}_{\Gamma_{0}}$ denotes a fundamental domain for the quotient space $X_{\Gamma_{0}}=\Gamma_{0}\backslash\mathbb{H}$.
+3. $L^{\infty}$-norm bounds for cusp forms revisited
 Refining arguments of [AM17] and [AM18], we first derive bounds for the Bergman kernel along the diagonal.
@@
 With notations as above, let $\Gamma\subset\mathrm{SL_{2}(\mathbb{R})}$ be a cofinite Fuchsian subgroup and $k\in\mathbb{R}_{\geq 5}$.
-Then, if $\Gamma$ is cocompact without elliptic elements, we have the bound
-(29) $\sup_{\tau\in\mathbb{H}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}=O_{\Gamma}(k).$
-Moreover, if $\Gamma$ is cofinite, we have the bound
-(30) $\sup_{\tau\in\mathbb{H}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}=O_{\Gamma}\big{(}k^{\frac{3}{2}}\big{)}.$
-The implied constants in the bounds (29) and (30) depend only on the Fuchsian subgroup $\Gamma$.
+Then, if $\Gamma$ is cocompact without elliptic elements, we have the $L^{\infty}$-norm bound
+(27) $\|B_{k,\chi}\|_{L^{\infty}}=\sup_{\tau\in\mathbb{H}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}=O_{\Gamma}(k).$
+Moreover, if $\Gamma$ is cofinite, we have the $L^{\infty}$-norm bound
+(28) $\|B_{k,\chi}\|_{L^{\infty}}=\sup_{\tau\in\mathbb{H}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}=O_{\Gamma}\big{(}k^{\frac{3}{2}}\big{)}.$
+The implied constants in the $L^{\infty}$-norm bounds (27) and (28) depend only on the Fuchsian subgroup $\Gamma$.
 Proof.
-When $\Gamma$ is cocompact without elliptic elements, the claimed bound (29) follows directly from Proposition 3.1.
+When $\Gamma$ is cocompact without elliptic elements, the claimed $L^{\infty}$-norm bound (27) follows directly from Proposition 3.1.
 Let next $\Gamma$ be a cofinite Fuchsian subgroup. From the proof of Theorem 6.1 in [FJK16], it follows that
-(31) $\sup_{\tau\in\mathbb{H}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}=\sup_{\begin{subarray}{c}\tau\in\partial\mathcal{F}_{Y}\\ Y=k/(2\pi)\end{subarray}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}},$
+(29) $\sup_{\tau\in\mathbb{H}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}=\sup_{\begin{subarray}{c}\tau\in\partial\mathcal{F}_{Y}\\ Y=k/(4\pi)\end{subarray}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}},$
 where $\partial\mathcal{F}_{Y}$ denotes the boundary of the truncated fundamental domain $\mathcal{F}_{Y}$ defined in (7). Combining Proposition 3.2 with (29) and employing the fact that
@@
 With notations as above, let $\Gamma\subset\mathrm{SL_{2}(\mathbb{R})}$ be a cofinite Fuchsian subgroup.
-For $k\in\mathbb{R}_{\geq 5}$, let $f\in S_{k,\chi}(\Gamma)$ be a cusp form, which is normalized with respect to the Petersson inner product (9).
-If $\Gamma$ is cocompact without elliptic elements, we have the bound
-(32) $\sup_{\tau\in\mathbb{H}}\|f(\tau)\|_{\mathrm{Pet}}^{2}=O_{\Gamma}(k).$
-Moreover, if $\Gamma$ is cofinite, we have the bound
-(33) $\sup_{\tau\in\mathbb{H}}\|f(\tau)\|_{\mathrm{Pet}}^{2}=O_{\Gamma}\big{(}k^{\frac{3}{2}}\big{)}.$
-The implied constants in the bounds (32) and (33) depend only on the Fuchsian subgroup $\Gamma$.
+For $k\in\mathbb{R}_{\geq 5}$, let $f\in S_{k,\chi}(\Gamma)$ be an $L^{2}$-normalized cusp form.
+If $\Gamma$ is cocompact without elliptic elements, we have the $L^{\infty}$-norm bound
+(30) $\|f\|_{L^{\infty}}=\sup_{\tau\in\mathbb{H}}\|f(\tau)\|_{\mathrm{Pet}}^{2}=O_{\Gamma}(k).$
+Moreover, if $\Gamma$ is cofinite, we have the $L^{\infty}$-norm bound
+(31) $\|f\|_{L^{\infty}}=\sup_{\tau\in\mathbb{H}}\|f(\tau)\|_{\mathrm{Pet}}^{2}=O_{\Gamma}\big{(}k^{\frac{3}{2}}\big{)}.$
+The implied constants in the $L^{\infty}$-norm bounds (30) and (31) depend only on the Fuchsian subgroup $\Gamma$.
 Proof.
@@
 With notations as above, let $\Gamma_{0}\subset\mathrm{SL_{2}}(\mathbb{R})$ be a fixed cofinite Fuchsian subgroup and let $\Gamma\subseteq\Gamma_{0}$ be a finite index subgroup of $\Gamma_{0}$.
-For $k\in\mathbb{R}_{\geq 5}$, let $f\in S_{k,\chi}(\Gamma)$ be a cusp form, which is normalized with respect to the Petersson inner product (9).
-If $\Gamma_{0}$ is cocompact without elliptic elements, we have the bound
-(34) $\sup_{\tau\in\mathbb{H}}\|f(\tau)\|_{\mathrm{Pet}}^{2}=O_{\Gamma_{0}}(k).$
-Moreover, if $\Gamma_{0}$ is cofinite, we have the bound
-(35) $\sup_{\tau\in\mathbb{H}}\|f(\tau)\|_{\mathrm{Pet}}^{2}=O_{\Gamma_{0}}\big{(}k^{\frac{3}{2}}\big{)}.$
-The implied constants in the bounds (34) and (35) depend only on the Fuchsian subgroup $\Gamma_{0}$.
+For $k\in\mathbb{R}_{\geq 5}$, let $f\in S_{k,\chi}(\Gamma)$ be an $L^{2}$-normalized cusp form.
+If $\Gamma_{0}$ is cocompact without elliptic elements, we have the $L^{\infty}$-norm bound
+(32) $\|f\|_{L^{\infty}}=\sup_{\tau\in\mathbb{H}}\|f(\tau)\|_{\mathrm{Pet}}^{2}=O_{\Gamma_{0}}(k).$
+Moreover, if $\Gamma_{0}$ is cofinite, we have the $L^{\infty}$-norm bound
+(33) $\|f\|_{L^{\infty}}=\sup_{\tau\in\mathbb{H}}\|f(\tau)\|_{\mathrm{Pet}}^{2}=O_{\Gamma_{0}}\big{(}k^{\frac{3}{2}}\big{)}.$
+The implied constants in the $L^{\infty}$-norm bounds (32) and (33) depend only on the Fuchsian subgroup $\Gamma_{0}$.
 Proof.
@@
 We are left to provide bounds for the pointwise Petersson norm of $f$, when $\tau$ ranges across the neighborhoods $\mathcal{F}^{Y}_{j}$ of the cusps.
-For this, we will have to distinguish between the two cases $Y>\frac{k}{2\pi}$ and $Y<\frac{k}{2\pi}$.
+For this, we will have to distinguish between the two cases $Y>\frac{k}{4\pi}$ and $Y<\frac{k}{4\pi}$.
 Without loss of generality, we can assume that $j=1$, when $p_{1}\in\mathcal{P}_{\Gamma}$ is the cusp at infinity for $\Gamma$ lying above the cusp $p$ at infinity for $\Gamma_{0}$ with ramification index $[\Gamma_{0,p}:\Gamma_{p_{1}}]$ (note that $\Gamma_{0,p}$ denotes the stabilizer subgroup of $p$ in $\Gamma_{0}$).
 With the above notations, we obtain the inclusion $\Gamma\setminus\Gamma_{p_{1}}\subseteq\Gamma_{0}\setminus\Gamma_{0,p}$.
-We first treat the case $Y>\frac{k}{2\pi}$, which implies that $\mathcal{F}_{1}^{Y}\subset\mathcal{F}_{1}^{k/(2\pi)}$.
+We first treat the case $Y>\frac{k}{4\pi}$, which implies that $\mathcal{F}_{1}^{Y}\subset\mathcal{F}_{1}^{k/(4\pi)}$.
 Arguing as in the proof of Theorem 6.1 in [FJK16], we deduce, recalling the inclusion $\Gamma\setminus\Gamma_{p_{1}}\subseteq\Gamma_{0}\setminus\Gamma_{0,p}$, that
-$\quad\sup_{\tau\in\mathcal{F}_{1}^{Y}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}\leq\sup_{\tau\in\mathcal{F}_{1}^{k/(2\pi)}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}\leq\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(2\pi)}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}$
-$\leq\frac{k-1}{4\pi}\Bigg{(}\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(2\pi)}}\sum_{\gamma\in\Gamma\setminus\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}+\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(2\pi)}}\sum_{\gamma\in\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}\Bigg{)}$
-$\leq\frac{k-1}{4\pi}\Bigg{(}\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(2\pi)}}\sum_{\gamma\in\Gamma_{0}\setminus\Gamma_{0,p}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}+\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(2\pi)}}\sum_{\gamma\in\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}\Bigg{)}.$
+$\quad\sup_{\tau\in\mathcal{F}_{1}^{Y}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}\leq\sup_{\tau\in\mathcal{F}_{1}^{k/(4\pi)}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}\leq\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(4\pi)}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}$
+$\leq\frac{k-1}{4\pi}\Bigg{(}\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(4\pi)}}\sum_{\gamma\in\Gamma\setminus\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}+\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(4\pi)}}\sum_{\gamma\in\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}\Bigg{)}$
+$\leq\frac{k-1}{4\pi}\Bigg{(}\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(4\pi)}}\sum_{\gamma\in\Gamma_{0}\setminus\Gamma_{0,p}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}+\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(4\pi)}}\sum_{\gamma\in\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}\Bigg{)}.$
 Arguments similar to the ones used to derive the bound (34), lead for the first term of (LABEL:thm6-eqn4) to the bound
-$\frac{k-1}{4\pi}\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(2\pi)}}\sum_{\gamma\in\Gamma_{0}\setminus\Gamma_{0,p}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}$
-$\quad\leq\frac{k-1}{2\pi}+\frac{3(k-1)}{\pi\cosh^{k-4}(r_{\Gamma_{0},k/(2\pi)}/4)}\bigg{(}1+\frac{1}{\sinh^{2}(r_{\Gamma_{0},k/(2\pi)}/4)}\bigg{)}+\frac{k-1}{4\pi}\sum_{e_{j}\in\mathcal{E}_{\Gamma_{0}}}(m_{e_{j}}-1),$
+$\frac{k-1}{4\pi}\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(4\pi)}}\sum_{\gamma\in\Gamma_{0}\setminus\Gamma_{0,p}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}$
+$\quad\leq\frac{k-1}{2\pi}+\frac{3(k-1)}{\pi\cosh^{k-4}(r_{\Gamma_{0},k/(4\pi)}/4)}\bigg{(}1+\frac{1}{\sinh^{2}(r_{\Gamma_{0},k/(4\pi)}/4)}\bigg{)}+\frac{k-1}{4\pi}\sum_{e_{j}\in\mathcal{E}_{\Gamma_{0}}}(m_{e_{j}}-1),$
 where
-$r_{\Gamma_{0},k/(2\pi)}=\inf\bigg{\{}\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)\,\bigg{|}\,\tau\in\partial\mathcal{F}_{1}^{k/(2\pi)},\,\gamma\in\Gamma_{0}\setminus\bigg{(}\Gamma_{0,p}\cup\bigcup_{e_{j}\in\mathcal{E}_{\Gamma_{0}}}\Gamma_{0,e_{j}}\bigg{)}\bigg{\}}>0.$
+$r_{\Gamma_{0},k/(4\pi)}=\inf\bigg{\{}\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)\,\bigg{|}\,\tau\in\partial\mathcal{F}_{1}^{k/(4\pi)},\,\gamma\in\Gamma_{0}\setminus\bigg{(}\Gamma_{0,p}\cup\bigcup_{e_{j}\in\mathcal{E}_{\Gamma_{0}}}\Gamma_{0,e_{j}}\bigg{)}\bigg{\}}>0.$
 Since it is easy to see that
-$\frac{1}{\sinh^{2}(r_{\Gamma_{0},k/(2\pi)})}=O_{\Gamma_{0}}(1),$
+$\frac{1}{\sinh^{2}(r_{\Gamma_{0},k/(4\pi)})}=O_{\Gamma_{0}}(1),$
 we arrive for the first term of (LABEL:thm6-eqn4) at the bound
-(39) $\frac{k-1}{4\pi}\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(2\pi)}}\sum_{\gamma\in\Gamma_{0}\setminus\Gamma_{0,p}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}=O_{\Gamma_{0}}(k).$
+(37) $\frac{k-1}{4\pi}\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(4\pi)}}\sum_{\gamma\in\Gamma_{0}\setminus\Gamma_{0,p}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}=O_{\Gamma_{0}}(k).$
 Using the same arguments as in the proof of Proposition 3.2, we derive for the second term of (LABEL:thm6-eqn4) the bound
-(40) $\frac{k-1}{4\pi}\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(2\pi)}}\sum_{\gamma\in\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}=O\big{(}k^{\frac{3}{2}}\big{)}.$
-By means of (LABEL:thm6-eqn4), we thus deduce from (39) and (40) in the case $Y>\frac{k}{2\pi}$ the bound
+(38) $\frac{k-1}{4\pi}\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(4\pi)}}\sum_{\gamma\in\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}=O\big{(}k^{\frac{3}{2}}\big{)}.$
+By means of (LABEL:thm6-eqn4), we thus deduce from (37) and (38) in the case $Y>\frac{k}{4\pi}$ the bound
 (39) $\sup_{\tau\in\mathcal{F}_{1}^{Y}}\|f(\tau)\|_{\mathrm{Pet}}^{2}\leq\sup_{\tau\in\mathcal{F}_{1}^{Y}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}=O_{\Gamma_{0}}\big{(}k^{\frac{3}{2}}\big{)}.$
-We finally turn to the case $Y<\frac{k}{2\pi}$, which implies that $\mathcal{F}_{1}^{k/(2\pi)}\subset\mathcal{F}_{1}^{Y}$. Here we find, arguing as in the preceding case that
-$\sup_{\tau\in\mathcal{F}_{1}^{Y}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}$ $\leq\sup_{\tau\in\mathcal{F}_{1}^{Y}\setminus\mathcal{F}_{1}^{k/(2\pi)}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}$
-$\leq\frac{k-1}{4\pi}\sup_{\tau\in\mathcal{F}_{1}^{Y}\setminus\mathcal{F}_{1}^{k/(2\pi)}}\sum_{\gamma\in\Gamma_{0}\setminus\Gamma_{0,p}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}$
-(42) $\quad\,+\frac{k-1}{4\pi}\sup_{\tau\in\mathcal{F}_{1}^{Y}\setminus\mathcal{F}_{1}^{k/(2\pi)}}\sum_{\gamma\in\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}.$
+We finally turn to the case $Y<\frac{k}{4\pi}$, which implies that $\mathcal{F}_{1}^{k/(4\pi)}\subset\mathcal{F}_{1}^{Y}$.
+Again, arguing as in the proof of Theorem 6.1 in [FJK16], we deduce after recalling once again the inclusion $\Gamma\setminus\Gamma_{p_{1}}\subseteq\Gamma_{0}\setminus\Gamma_{0,p}$ that
+$\sup_{\tau\in\mathcal{F}_{1}^{Y}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}$ $\leq\sup_{\tau\in\mathcal{F}_{1}^{Y}\setminus\mathcal{F}_{1}^{k/(4\pi)}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}$
+$\leq\frac{k-1}{4\pi}\sup_{\tau\in\mathcal{F}_{1}^{Y}\setminus\mathcal{F}_{1}^{k/(4\pi)}}\sum_{\gamma\in\Gamma_{0}\setminus\Gamma_{0,p}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}$
+(40) $\quad\,+\frac{k-1}{4\pi}\sup_{\tau\in\mathcal{F}_{1}^{Y}\setminus\mathcal{F}_{1}^{k/(4\pi)}}\sum_{\gamma\in\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}.$
 As before, we now obtain the bounds
-$\frac{k-1}{4\pi}\sup_{\tau\in\mathcal{F}_{1}^{Y}\setminus\mathcal{F}_{1}^{k/(2\pi)}}\sum_{\gamma\in\Gamma_{0}\setminus\Gamma_{0,p}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}$
+$\frac{k-1}{4\pi}\sup_{\tau\in\mathcal{F}_{1}^{Y}\setminus\mathcal{F}_{1}^{k/(4\pi)}}\sum_{\gamma\in\Gamma_{0}\setminus\Gamma_{0,p}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}$
 $\quad\leq\frac{k-1}{2\pi}+\frac{3(k-1)}{\pi\cosh^{k-4}(r^{\prime}_{\Gamma_{0},Y}/4)}\bigg{(}1+\frac{1}{\sinh^{2}(r^{\prime}_{\Gamma_{0},Y}/4)}\bigg{)}+\frac{k-1}{4\pi}\sum_{e_{j}\in\mathcal{E}_{\Gamma_{0}}}(m_{e_{j}}-1)$
@@
 as well as the bound
-(44) $\frac{k-1}{4\pi}\sup_{\tau\in\mathcal{F}_{1}^{Y}\setminus\mathcal{F}_{1}^{k/(2\pi)}}\sum_{\gamma\in\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}=O\big{(}k^{\frac{3}{2}}\big{)}.$
-By means of (42), we thus deduce from (43) and (44) in the case $Y<\frac{k}{2\pi}$ the bound
+(42) $\frac{k-1}{4\pi}\sup_{\tau\in\mathcal{F}_{1}^{Y}\setminus\mathcal{F}_{1}^{k/(4\pi)}}\sum_{\gamma\in\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}=O\big{(}k^{\frac{3}{2}}\big{)}.$
+By means of (40), we thus deduce from (41) and (42) in the case $Y<\frac{k}{4\pi}$ the bound
 (43) $\sup_{\tau\in\mathcal{F}_{1}^{Y}}\|f(\tau)\|_{\mathrm{Pet}}^{2}\leq\sup_{\tau\in\mathcal{F}_{1}^{Y}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}=O_{\Gamma_{0},Y}\big{(}k^{\frac{3}{2}}\big{)}.$
@@
 If $f\in S_{k,\chi}(\Gamma)$ is not a Hecke eigenform, then there is no evidence from the literature to suggest that the estimates (30) and (31) can be improved.
-Thus the estimates (32) and (33) are expected to be optimal.
+Thus, the estimates (30) and (31) are expected to be optimal.
 4. Sup-norm bounds for Jacobi cusp forms
-For $k\in\mathbb{Z}_{\geq 5}$ and $m\in\mathbb{Z}_{\geq 1}$, let $f\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$ be a Jacobi cusp form of weight $k$ and index $m$ for the full modular group $\Gamma_{0}=\mathrm{SL}_{2}(\mathbb{Z})$, which is normalized with respect to the Petersson inner product defined by (12).
-We now aim at bounding the quantity
-$\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|f(\tau,z)\|_{\mathrm{Pet}}.$
-For this we will employ the decomposition (13), namely
-$f(\tau,z)=\sum_{\mu=0}^{2m-1}\varphi_{\mu}(\tau)\vartheta_{\mu,m}(\tau,z),$
-where the functions $\varphi_{\mu}$ are cusp forms of weight $(k-\frac{1}{2})$ with respect to the finite index subgroup $\Gamma_{1}=\Gamma_{0}(4m)$ of $\Gamma_{0}$ and the theta functions $\vartheta_{\mu,m}$ are defined in (14).
-For a fixed $\tau=\xi+i\eta\in\mathbb{H}$, consider the elliptic curve $E_{\tau}\coloneqq\mathbb{C}/\Lambda_{\tau}$ with $\Lambda_{\tau}\coloneqq\mathbb{Z}\oplus\,\tau\mathbb{Z}$.
-Let $O_{\tau}$ denote the identity element of $E_{\tau}$, when considered as an abelian group with $\oplus_{\tau}$ denoting the group operation, and let
-$[2]\colon E_{\tau}\longrightarrow E_{\tau}$
-be multiplication by $2$, given by the assignment $z\mapsto 2z\coloneqq z\oplus_{\tau}z$, which is an isogeny of degree $4$.
-Let $\mathcal{M}_{\tau}$ be the line bundle associated to the divisor $O_{\tau}$.
-Then, the theorem of the cube gives the isomorphism
-$[2]^{\ast}\mathcal{M}_{\tau}\cong\mathcal{M}_{\tau}^{\otimes{4}}.$
-We then put $\mathcal{L}_{\tau}\coloneqq\mathcal{M}_{\tau}^{\otimes{2}}$ and find that the theta functions $\vartheta_{\mu,m}(\tau,\cdot)$ ($\mu=0,\ldots,2m-1$) arise as global holomorphic sections of the line bundle $\mathcal{L}_{\tau}^{\otimes{m}}$.
-The pointwise norm of $\vartheta_{\mu,m}(\tau,\cdot)\in H^{0}(E_{\tau},\mathcal{L}_{\tau}^{\otimes{m}})$ at the point $z=x+iy\in E_{\tau}$ (identifying $E_{\tau}$ with its universal cover $\mathbb{C}$) is given by the following formula
-(46) $\|\vartheta_{\mu,m}(\tau,z)\|_{\mathcal{L}_{\tau}^{\otimes{m}}}^{2}\coloneqq|\vartheta_{\mu,m}(\tau,z)|^{2}\,\sqrt{\eta}\,e^{-\frac{4\pi my^{2}}{\eta}}.$
-Let $\mu_{\mathrm{eucl}}$ denote the Euclidean metric on $E_{\tau}$; at the point $z=x+iy\in E_{\tau}$, it is given by the formula
-(47) $\mu_{\mathrm{eucl}}(z)=\frac{i}{2}\cdot\frac{\mathrm{d}z\wedge\mathrm{d}\overline{z}}{\eta}=\frac{\mathrm{d}x\wedge\mathrm{d}y}{\eta}.$
-Furthermore, the $\mathrm{L}^{2}$-inner product of $\vartheta_{\mu,m}(\tau,\cdot),\vartheta_{\mu^{\prime},m}(\tau,\cdot)\in H^{0}(E_{\tau},\mathcal{L}_{\tau}^{\otimes{m}})$ is given by the formula
-$\langle\vartheta_{\mu,m}(\tau,\cdot),\vartheta_{\mu^{\prime},m}(\tau,\cdot)\rangle_{\mathrm{L}^{2},\mathcal{L}_{\tau}^{\otimes{m}}}$ $\coloneqq\int\limits_{E_{\tau}}\vartheta_{\mu,m}(\tau,z)\,\overline{\vartheta_{\mu^{\prime},m}(\tau,z)}\,\sqrt{\eta}\,e^{-\frac{4\pi my^{2}}{\eta}}\,\mu_{\mathrm{eucl}}(z)$
-(48) $=\int\limits_{0}^{\eta}\int\limits_{0}^{1}\vartheta_{\mu,m}(\tau,z)\,\overline{\vartheta_{\mu^{\prime},m}(\tau,z)}\,\sqrt{\eta}\,e^{-\frac{4\pi my^{2}}{\eta}}\,\frac{\mathrm{d}x\wedge\mathrm{d}y}{\eta}.$
-The $\mathrm{L}^{2}$-norm of $\vartheta_{\mu,m}(\tau,\cdot)\in H^{0}(E_{\tau},\mathcal{L}_{\tau}^{\otimes{m}})$ induced by the $\mathrm{L}^{2}$-inner product defined above is thus given by
-(49) $\|\vartheta_{\mu,m}(\tau,\cdot)\|_{\mathrm{L}^{2},\,\mathcal{L}_{\tau}^{\otimes{m}}}^{2}=\langle\vartheta_{\mu,m}(\tau,\cdot),\vartheta_{\mu,m}(\tau,\cdot)\rangle_{\mathrm{L}^{2},\mathcal{L}_{\tau}^{\otimes{m}}}.$
-We next recall the well-known fact about the orthogonality of the Jacobi theta functions (which was already implicitly used in formula (15)).
+For $k\in\mathbb{Z}_{\geq 5}$ and $m\in\mathbb{Z}_{\geq 1}$, let $\phi\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$ be an $L^{2}$-normalized Jacobi cusp form of weight $k$ and index $m$ for the full modular group $\Gamma_{0}=\mathrm{SL}_{2}(\mathbb{Z})$.
+We next aim at bounding the quantity
+$\|\phi\|_{L^{\infty}}=\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|\phi(\tau,z)\|_{\mathrm{Pet}}.$
+By the $\Gamma_{0}\ltimes\mathbb{Z}^{2}$-invariance of the function $\|\phi(\tau,z)\|_{\mathrm{Pet}}$ on $\mathbb{H}\times\mathbb{C}$, it suffices to find a bound for the quantity $\|\phi(\tau,z)\|_{\mathrm{Pet}}$, when $\tau$ is ranging through a fundamental domain $\mathcal{F}_{\Gamma_{0}}$ for $\Gamma_{0}$ and $z$ is ranging through a fundamental domain of the elliptic curve $\mathbb{C}/(\mathbb{Z}\oplus\tau\mathbb{Z})$.
+Without loss of generality, we can assume in the sequel that
+$\mathcal{F}_{\Gamma_{0}}=\{\tau=\xi+i\eta\in\mathbb{H}\,|\,-1/2\leq\xi\leq 1/2,\,|\tau|\geq 1\},$
+and, given $\tau=\xi+i\eta\in\mathcal{F}_{\Gamma_{0}}$, that
+[…]
+Furthermore, in order to arrive at an $L^{\infty}$-bound of the quantity $\|\phi(\tau,z)\|_{\mathrm{Pet}}$, we will make use of the decomposition (13), namely
+$\phi(\tau,z)=\sum_{\mu=0}^{2m-1}h_{\mu}(\tau)\vartheta_{\mu,m}(\tau,z),$
+where the functions $h_{\mu}$ are cusp forms of weight $(k-\frac{1}{2})$ with respect to the finite index subgroup $\Gamma_{1}=\Gamma_{0,1}(4m)$ of $\Gamma_{0}$ and the theta functions $\vartheta_{\mu,m}$ are defined in (14).
 Lemma 4.1.
-The set of Jacobi theta functions $\{\vartheta_{\mu,m}(\tau,\cdot)\}_{\mu=0}^{2m-1}$ constitutes an orthogonal basis for the space $H^{0}(E_{\tau},\mathcal{L}_{\tau}^{\otimes{m}})$ with respect to the $\mathrm{L}^{2}$-inner product (48).
+With notations as above, we have the bound
+(44) $\sum_{\mu=0}^{2m-1}\|\vartheta_{\mu,m}(\tau,z)\|^{2}_{\mathrm{Pet}}\leq 2m\,\mathrm{Im}(\tau)^{\frac{1}{2}}+O\big{(}m^{\frac{1}{2}}\big{)}$
+for $(\tau,z)\in\mathbb{H}\times\mathbb{C}$.
 Proof.
-As above, we write $\tau=\xi+i\eta\in\mathbb{H}$ and for $\mu,\mu^{\prime}=0,\ldots,2m-1$ and $n\in\mathbb{Z}$, we define
-(50) $a(\mu,n)\coloneqq n-\frac{\mu}{2m}\quad\text{and}\quad b(\mu,n)\coloneqq 2mn-\mu.$
-Recalling the definition of the Jacobi theta function (14), we compute for the $\mathrm{L}^{2}$-inner product (48)
-$\langle\vartheta_{\mu,m}(\tau,\cdot),\vartheta_{\mu^{\prime},m}(\tau,\cdot)\rangle_{\mathrm{L}^{2},\,\mathcal{L}_{\tau}^{\otimes{m}}}$
-(51) $\qquad=\int\limits_{0}^{\eta}\sum_{n\in\mathbb{Z}}\,\sum_{n^{\prime}\in\mathbb{Z}}e^{2\pi im\tau a(\mu,n)^{2}}\overline{e^{2\pi im\tau a(\mu^{\prime},n^{\prime})^{2}}}\bigg{(}\int\limits_{0}^{1}e^{2\pi izb(\mu,n)}\overline{e^{2\pi izb(\mu^{\prime},n^{\prime})}}\,\mathrm{d}x\bigg{)}\,e^{-\frac{4\pi my^{2}}{\eta}}\,\frac{\mathrm{d}y}{\sqrt{\eta}}.$
-Observing that the inequalities $-2m<\mu-\mu^{\prime}<2m$ imply that the equality
-$b(\mu,n)=b(\mu^{\prime},n^{\prime})\quad\Longleftrightarrow\quad 2m(n-n^{\prime})=\mu-\mu^{\prime}$
-can only hold if $n=n^{\prime}$ and $\mu=\mu^{\prime}$, we deduce
-$\int\limits_{0}^{1}e^{2\pi izb(\mu,n)}\overline{e^{2\pi izb(\mu^{\prime},n^{\prime})}}\,\mathrm{d}x$ $=e^{-2\pi y(b(\mu,n)+b(\mu^{\prime},n^{\prime}))}\int\limits_{0}^{1}e^{2\pi ix(b(\mu,n)-b(\mu^{\prime},n^{\prime}))}\,\mathrm{d}x$
-(52) =$=\begin{cases}e^{-4\pi yb(\mu,n)},&\text{if }\mu=\mu^{\prime}\text{ and }n=n^{\prime},\\ 0,&\text{else}.\end{cases}$
-This proves the pairwise orthogonality of the $2m$ Jacobi theta functions $\{\vartheta_{\mu,m}(\tau,\cdot)\}_{\mu=0}^{2m-1}$ and thus their linear independence.
-Now, since we have $\deg(\mathcal{L}_{\tau}^{\otimes{m}})=2m$, the Riemann–Roch Theorem on the elliptic curve $E_{\tau}$ shows that
-$\dim H^{0}(E_{\tau},\mathcal{L}_{\tau}^{\otimes{m}})=2m,$
-from which the claim follows. ∎
-Proposition 4.2.
-With notations as above, we have for $\mu=0,\ldots,2m-1$, the bound
-$\sup_{\tau\in\mathbb{H}}\|\vartheta_{\mu,m}(\tau,\cdot)\|_{\mathrm{L}^{2},\,\mathcal{L}_{\tau}^{\otimes{m}}}^{2}=O\bigg{(}\frac{1}{\sqrt{m}}\bigg{)}.$
+By definition, we have for $(\tau,z)\in\mathbb{H}\times\mathbb{C}$ the equalities
+$\vartheta_{\mu,m}(\tau,z)$ $=\sum_{n\in\mathbb{Z}}e^{2\pi im\tau(n-\frac{\mu}{2m})^{2}+2\pi iz(2mn-\mu)}$
+$=\sum_{n\in\mathbb{Z}}e^{2\pi imn^{2}\tau-2\pi i\cdot 2mn\cdot\frac{\mu\tau}{2m}+\pi i\frac{\mu^{2}\tau}{2m}+2\pi i\cdot 2mn\cdot z-2\pi i\mu z}$
+$=e^{\pi i\frac{\mu^{2}\tau}{2m}-2\pi i\mu z}\,\sum_{n\in\mathbb{Z}}e^{2\pi imn^{2}\tau+2\pi i(z-\frac{\mu\tau}{2m})2mn}$
+$=\vartheta_{0,m}\Big{(}\tau,z-\frac{\mu\tau}{2m}\Big{)}\,e^{\pi i\frac{\mu^{2}\tau}{2m}-2\pi i\mu z}.$
+Taking absolute values, we get
+$|\vartheta_{\mu,m}(\tau,z)|=\Big{|}\vartheta_{0,m}\Big{(}\tau,z-\frac{\mu\tau}{2m}\Big{)}\Big{|}\,e^{-\pi\frac{\mu^{2}\eta}{2m}+2\pi\mu y}.$
+This implies the relation
+$\|\vartheta_{\mu,m}(\tau,z)\|_{\mathrm{Pet}}$ $=|\vartheta_{\mu,m}(\tau,z)|\,\eta^{\frac{1}{4}}\,e^{-\frac{2\pi my^{2}}{\eta}}$
+$=\Big{|}\vartheta_{0,m}\Big{(}\tau,z-\frac{\mu\tau}{2m}\Big{)}\Big{|}\,\eta^{\frac{1}{4}}\,e^{-\pi\frac{\mu^{2}\eta}{2m}+2\pi\mu y-2\pi m\frac{y^{2}}{\eta}}$
+$=\Big{|}\vartheta_{0,m}\Big{(}\tau,z-\frac{\mu\tau}{2m}\Big{)}\Big{|}\,\eta^{\frac{1}{4}}\,e^{-\frac{2\pi m}{\eta}(y-\frac{\mu\eta}{2m})^{2}}.$
+From this, we derive for $(\tau,z)\in\mathbb{H}\times\mathbb{C}$ the bound
+$\|\vartheta_{\mu,m}(\tau,z)\|_{\mathrm{Pet}}$ $=\Big{|}\vartheta_{0,m}\Big{(}\tau,z-\frac{\mu\tau}{2m}\Big{)}\Big{|}\,\eta^{\frac{1}{4}}\,e^{-\frac{2\pi m}{\eta}(y-\frac{\mu\eta}{2m})^{2}}$
+$\leq\sum_{n\in\mathbb{Z}}e^{-2\pi mn^{2}\eta-4\pi mn(y-\frac{\mu\eta}{2m})-\frac{2\pi m}{\eta}(y-\frac{\mu\eta}{2m})^{2}}\,\eta^{\frac{1}{4}}$
+$=\sum_{n\in\mathbb{Z}}e^{-\frac{2\pi m}{\eta}(n\eta+y-\frac{\mu\eta}{2m})^{2}}\,\eta^{\frac{1}{4}}.$
+By an integral test, we arrive for $(\tau,z)\in\mathbb{H}\times\mathbb{C}$ at the bound
+$\|\vartheta_{\mu,m}(\tau,z)\|_{\mathrm{Pet}}$ $\leq\bigg{(}1+\int_{-\infty}^{\infty}e^{-\frac{2\pi m}{\eta}(n\eta+y-\frac{\mu\eta}{2m})^{2}}\,\mathrm{d}n\bigg{)}\eta^{\frac{1}{4}}$
+$=\bigg{(}1+\frac{1}{\sqrt{2m\eta}}\bigg{)}\eta^{\frac{1}{4}},$
+which leads to
+$\sum_{\mu=0}^{2m-1}\|\vartheta_{\mu,m}(\tau,z)\|^{2}_{\mathrm{Pet}}\leq 2m\eta^{\frac{1}{2}}+O\big{(}m^{\frac{1}{2}}\big{)}.$
+This completes the proof of the lemma. ∎
+Lemma 4.2.
+With notations as above, let $k\in\mathbb{Z}_{\geq 5}$, let $\Gamma\subseteq\Gamma_{0}$ be a subgroup of finite index, whose stabilizer subgroup $\Gamma_{i\infty}$ is generated by $\big{(}\begin{smallmatrix}1&1\\ 0&1\end{smallmatrix}\big{)}$, and let $f\in S_{k,\chi}(\Gamma)$ be an $L^{2}$-normalized cusp form of weight $k$ and character $\chi$ for $\Gamma$.
+Then, we have the bound
+(45) $\sup_{\tau\in\mathcal{F}_{\Gamma_{0}}}\big{(}\|f(\tau)\|_{\mathrm{Pet}}^{2}\,\mathrm{Im}(\tau)^{\frac{1}{2}}\big{)}=O_{\Gamma_{0}}(k^{2}),$
+where the implied constant depends on $\Gamma_{0}$.
 Proof.
-Let $\tau=\xi+i\eta\in\mathbb{H}$, $z=x+iy\in\mathbb{C}$, and $\mu=0,\ldots,2m-1$.
-For $\tau\in\mathbb{H}$, we need to bound the $\mathrm{L}^{2}$-norm (49) of $\vartheta_{\mu,m}(\tau,\cdot)\in H^{0}(E_{\tau},\mathcal{L}_{\tau}^{\otimes{m}})$, namely the quantity
-$\|\vartheta_{\mu,m}(\tau,\cdot)\|_{\mathrm{L}^{2},\,\mathcal{L}_{\tau}^{\otimes{m}}}^{2}=\langle\vartheta_{\mu,m}(\tau,\cdot),\vartheta_{\mu,m}(\tau,\cdot)\rangle_{\mathrm{L}^{2},\,\mathcal{L}_{\tau}^{\otimes{m}}}.$
-Recalling the notation (50) and introducing $a(n)\coloneqq a(\mu,n)$ as well as $b(n)\coloneqq b(\mu,n)$ as a short-hand, we arrive as in (51) at the equality
-$\|\vartheta_{\mu,m}(\tau,\cdot)\|_{\mathrm{L}^{2},\,\mathcal{L}_{\tau}^{\otimes{m}}}^{2}$
-(53) $\qquad=\int\limits_{0}^{\eta}\sum_{n\in\mathbb{Z}}\,\sum_{n^{\prime}\in\mathbb{Z}}e^{2\pi im\tau a(n)^{2}}\overline{e^{2\pi im\tau a(n^{\prime})^{2}}}\bigg{(}\int\limits_{0}^{1}e^{2\pi izb(n)}\overline{e^{2\pi izb(n^{\prime})}}\,\mathrm{d}x\bigg{)}\,e^{-\frac{4\pi my^{2}}{\eta}}\,\frac{\mathrm{d}y}{\sqrt{\eta}}.$
-Since (by (52))
-$\int\limits_{0}^{1}e^{2\pi izb(n)}\overline{e^{2\pi izb(n^{\prime})}}\,\mathrm{d}x=\begin{cases}e^{-4\pi yb(n)},&\text{ if }n=n^{\prime},\\ 0,&\text{ else},\end{cases}$
-we arrive at
-(54) $\sum_{n\in\mathbb{Z}}\,\sum_{n^{\prime}\in\mathbb{Z}}e^{2\pi im\tau a(n)^{2}}\overline{e^{2\pi im\tau a(n^{\prime})^{2}}}\int\limits_{0}^{1}e^{2\pi izb(n)}\overline{e^{2\pi izb(n^{\prime})}}\,\mathrm{d}x=\sum_{n\in\mathbb{Z}}e^{-4\pi m\eta\big{(}n-\frac{\mu}{2m}\big{)}^{2}-4\pi y(2mn-\mu)}.$
-Substituting (54) into (53) and using an integral test, we find the bound
-$\|\vartheta_{\mu,m}(\tau,\cdot)\|_{\mathrm{L}^{2},\,\mathcal{L}_{\tau}^{\otimes{m}}}^{2}$ $=\sum_{n\in\mathbb{Z}}\int\limits_{0}^{\eta}e^{-4\pi m\eta\big{(}n-\frac{\mu}{2m}\big{)}^{2}-4\pi y(2mn-\mu)}\,e^{-\frac{4\pi my^{2}}{\eta}}\,\frac{\mathrm{d}y}{\sqrt{\eta}}$
-(55) $\leq\int\limits_{0}^{\eta}\int\limits_{-\infty}^{\infty}e^{-4\pi m\eta\big{(}\nu-\frac{\mu}{2m}\big{)}^{2}-4\pi y(2m\nu-\mu)-\frac{4\pi my^{2}}{\eta}}\,\mathrm{d}\nu\,\frac{\mathrm{d}y}{\sqrt{\eta}}$
-(56) $\quad+\int\limits_{0}^{\eta}e^{-\frac{\pi\eta\mu^{2}}{m}+4\pi y\mu-\frac{4\pi my^{2}}{\eta}}\,\frac{\mathrm{d}y}{\sqrt{\eta}}.$
-Now, we rewrite the exponent of the integrand in (55) in the form
-$4\pi m\eta\bigg{(}\nu-\frac{\mu}{2m}\bigg{)}^{2}+4\pi y(2m\nu-\mu)+\frac{4\pi my^{2}}{\eta}$
-$\qquad=\bigg{(}\sqrt{\frac{\pi\eta}{m}}(2m\nu-\mu)\bigg{)}^{2}+4\pi y(2m\nu-\mu)+\bigg{(}2\sqrt{\frac{\pi m}{\eta}}y\bigg{)}^{2}$
-$\qquad=\bigg{(}\sqrt{\frac{\pi\eta}{m}}(2m\nu-\mu)+2\sqrt{\frac{\pi m}{\eta}}y\bigg{)}^{2}.$
-Substituting
-$\rho\coloneqq\sqrt{\frac{\pi\eta}{m}}(2m\nu-\mu)+2\sqrt{\frac{\pi m}{\eta}}y$
-into (55), we obtain for the inner integral
-$\int\limits_{-\infty}^{\infty}e^{-4\pi m\eta\big{(}\nu-\frac{\mu}{2m}\big{)}^{2}-4\pi y(2m\nu-\mu)-\frac{4\pi my^{2}}{\eta}}\,\mathrm{d}\nu=\frac{1}{2\sqrt{\pi m\eta}}\int\limits_{-\infty}^{\infty}e^{-\rho^{2}}\,\mathrm{d}\rho=\frac{1}{2\sqrt{m}}\frac{1}{\sqrt{\eta}}.$
-From this, we compute the double integral (55) as
-(57) $\int\limits_{0}^{\eta}\int\limits_{-\infty}^{\infty}e^{-4\pi m\eta\big{(}\nu-\frac{\mu}{2m}\big{)}^{2}-4\pi y(2m\nu-\mu)-\frac{4\pi my^{2}}{\eta}}\,\mathrm{d}\nu\,\frac{\mathrm{d}y}{\sqrt{\eta}}=\frac{1}{2\sqrt{m}}\frac{1}{\sqrt{\eta}}\int\limits_{0}^{\eta}\frac{\mathrm{d}y}{\sqrt{\eta}}=\frac{1}{2\sqrt{m}}.$
-For the integral (56), we find in a similar way
-(58) $\int\limits_{0}^{\eta}e^{-\frac{\pi\eta\mu^{2}}{m}+4\pi y\mu-\frac{4\pi my^{2}}{\eta}}\,\frac{\mathrm{d}y}{\sqrt{\eta}}\leq\int\limits_{-\infty}^{\infty}e^{-\big{(}-\sqrt{\frac{\pi\eta}{m}}\mu+2\sqrt{\frac{\pi m}{\eta}}y\big{)}^{2}}\,\frac{\mathrm{d}y}{\sqrt{\eta}}=\frac{1}{2\sqrt{m}}.$
-Adding up the bounds (57) and (58), yields
-$\|\vartheta_{\mu,m}(\tau,\cdot)\|_{\mathrm{L}^{2},\,\mathcal{L}_{\tau}^{\otimes{m}}}^{2}\leq\frac{1}{\sqrt{m}},$
-which proves the claim. ∎
-Proposition 4.3.
-With notations as above, we have the bound
-(59) $\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\sum_{\mu=0}^{2m-1}\|\vartheta_{\mu,m}(\tau,z)\|_{\mathcal{L}_{\tau}^{\otimes{m}}}^{2}=O(\sqrt{m}).$
+Since the width of the cusp $p_{1}=i\infty$ of $\Gamma$ equals $1$, the neighborhood $\mathcal{F}_{1}^{Y}$ of $p_{1}$ is contained in the fundamental domain $\mathcal{F}_{\Gamma_{0}}$, provided that $Y>1$.
+For a given $L^{2}$-normalized $f\in S_{k,\chi}(\Gamma)$, we then write
+$\|f(\tau)\|_{\mathrm{Pet}}^{2}\,\mathrm{Im}(\tau)^{\frac{1}{2}}=\bigg{|}\frac{f(\tau)}{e^{2\pi i\tau}}\bigg{|}^{2}\cdot\frac{\eta^{k+1/2}}{e^{4\pi\eta}}.$
+Now, the function $|f(\tau)/e^{2\pi i\tau}|^{2}$ is subharmonic and bounded in the neighborhood $\mathcal{F}_{1}^{Y}$ and thus takes its maximum on the boundary $\partial\mathcal{F}_{1}^{Y}$ by the maximum principle for subharmonic functions.
+Furthermore, by elementary calculus, we find that the function $\eta^{k+1/2}/e^{4\pi\eta}$ attains its maximum at
+$\eta_{0}=\frac{k+1/2}{4\pi},$
+and is strictly monotonically decreasing for $\eta>\eta_{0}$. By assuming for the moment that $k>12$ and by choosing
+$Y\coloneqq\eta_{0}=\frac{k+1/2}{4\pi}>1,$
+we derive from the arguments used in the proof of Theorem 3.5, in particular from the bound (LABEL:thm6-eqn4), that
+$\sup_{\tau\in\mathcal{F}_{\Gamma_{0}}}\big{(}\|f(\tau)\|_{\mathrm{Pet}}^{2}\,\mathrm{Im}(\tau)^{\frac{1}{2}}\big{)}=\sup_{\tau\in\mathcal{F}_{1}^{\eta_{0}}}\big{(}\|f(\tau)\|_{\mathrm{Pet}}^{2}\,\eta^{\frac{1}{2}}\big{)}$
+$\qquad\quad=\sup_{\tau\in\partial\mathcal{F}_{1}^{\eta_{0}}}\big{(}\|f(\tau)\|_{\mathrm{Pet}}^{2}\,\eta^{\frac{1}{2}}\big{)}\leq\sup_{\tau\in\partial\mathcal{F}_{1}^{\eta_{0}}}\big{(}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}\,\eta^{\frac{1}{2}}\big{)}$
+(46) $\qquad\quad\leq k^{\frac{3}{2}}\bigg{(}\sup_{\tau\in\partial\mathcal{F}_{1}^{\eta_{0}}}\sum_{\gamma\in\Gamma\setminus\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}+\sup_{\tau\in\partial\mathcal{F}_{1}^{\eta_{0}}}\sum_{\gamma\in\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}\bigg{)}.$
+Finally, again using the arguments given in the proof of Theorem 3.5, shows that
+(47) $\sup_{\tau\in\partial\mathcal{F}_{1}^{\eta_{0}}}\sum_{\gamma\in\Gamma\setminus\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}=O_{\Gamma_{0}}(1),$
+(48) $\sup_{\tau\in\partial\mathcal{F}_{1}^{\eta_{0}}}\sum_{\gamma\in\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}=O_{\Gamma_{0}}\big{(}k^{\frac{1}{2}}\big{)}.$
+Combining the estimate (46) with the bounds (47) and (48), establishes the proof of the lemma for $k>12$.
+The finitely many remaining cases $k\in\{5,\ldots,12\}$ can now be obviously added to the statement by possibly enlarging the implied constant.
+∎
+Lemma 4.3.
+For the index of the congruence subgroup
+$\Gamma_{1}=\Gamma_{0,1}(4m)=\bigg{\{}\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\Gamma_{0}\,\bigg{|}\,c\equiv 0,\,d\equiv 1\mod 4m\bigg{\}}$
+in $\Gamma_{0}=\mathrm{SL}_{2}(\mathbb{Z})$, we have
+$[\Gamma_{0}:\Gamma_{1}]=O_{\epsilon}\big{(}m^{2+\epsilon}\big{)},$
+where the implied constant depends on the choice of $\epsilon>0$.
 Proof.
-We start by considering the Bergman kernel $B_{\mathcal{L}_{\tau}^{\otimes{m}}}(z)$ associated to the line bundle $\mathcal{L}_{\tau}^{\otimes{m}}$ on the elliptic curve $E_{\tau}$ equipped with the Kähler form $\omega=\mu_{\mathrm{eucl}}$ (see formula (47)).
-From the bound (17), we recall
-$B_{\mathcal{L}_{\tau}^{\otimes{m}}}(z)=\mathrm{det}_{\omega}\big{(}\mathrm{c}_{1}(\mathcal{L}_{\tau},\|\,\cdot\,\|_{\mathcal{L}_{\tau}})(z)\big{)}\,m+O(1),$
-where $\mathrm{c}_{1}(\mathcal{L}_{\tau},\|\,\cdot\,\|_{\mathcal{L}_{\tau}})(z)$ is the curvature form of $\mathcal{L}_{\tau}$ at the point $z\in E_{\tau}$.
-Using equations (16) and (46), the curvature form $\mathrm{c}_{1}(\mathcal{L}_{\tau},\|\,\cdot\,\|_{\mathcal{L}_{\tau}})(z)$ is given by the formula (identifying again $E_{\tau}$ with its universal cover $\mathbb{C}$ and writing $z=x+iy$)
-$\mathrm{c}_{1}(\mathcal{L}_{\tau},\|\,\cdot\,\|_{\mathcal{L}_{\tau}})(z)$ $=-\frac{i}{2\pi}\partial_{z}\partial_{\overline{z}}\log\|\vartheta_{\mu,1}(\tau,z)\|_{\mathcal{L}_{\tau}}^{2}$
-$=-\frac{i}{2\pi}\partial_{z}\partial_{\overline{z}}\bigg{(}-\frac{4\pi y^{2}}{\eta}\bigg{)}=2\mu_{\mathrm{eucl}}(z)=2\omega(z).$
-This yields $\mathrm{det}_{\omega}\big{(}\mathrm{c}_{1}(\mathcal{L}_{\tau},\|\,\cdot\,\|_{\mathcal{L}_{\tau}})(z)\big{)}=2$, and therefore we get
-(60) $B_{\mathcal{L}_{\tau}^{\otimes{m}}}(z)=2m+O(1).$
-Setting
-$\Theta_{\mu,m}(\tau,z)\coloneqq\frac{\vartheta_{\mu,m}(\tau,z)}{\|\vartheta_{\mu,m}(\tau,\cdot)\|_{\mathrm{L}^{2},\mathcal{L}^{\otimes{m}}}}\qquad(\mu=0,\ldots,2m-1)$
-and using Lemma 4.1, we deduce that the set $\{\Theta_{\mu,m}(\tau,\cdot)\}_{\mu=0}^{2m-1}$ constitutes an orthonormal basis for the space $H^{0}(E_{\tau},\mathcal{L}_{\tau}^{\otimes{m}})$ with respect to the $\mathrm{L}^{2}$-inner product (48).
-By the very definition of the Bergman kernel we thus have
-$B_{\mathcal{L}_{\tau}^{\otimes{m}}}(z)=\sum_{\mu=0}^{2m-1}\|\Theta_{\mu,m}(\tau,z)\|_{\mathcal{L}_{\tau}^{\otimes{m}}}^{2}$
-for any $z\in E_{\tau}$. Finally, using Proposition 4.2 and observing that
-$\sum_{\mu=0}^{2m-1}\|\vartheta_{\mu,m}(\tau,z)\|_{\mathcal{L}_{\tau}^{\otimes{m}}}^{2}\leq\frac{1}{\sqrt{m}}\sum_{\mu=0}^{2m-1}\|\Theta_{\mu,m}(\tau,z)\|_{\mathcal{L}_{\tau}^{\otimes{m}}}^{2}=\frac{1}{\sqrt{m}}B_{\mathcal{L}_{\tau}^{\otimes{m}}}(z),$
-the claim follows by means of (60). ∎
+The group $\Gamma_{0,1}(4m)$ fits into the short exact sequence
+$0\longrightarrow\Gamma_{0,1}(4m)\longrightarrow\Gamma_{0}(4m)\longrightarrow(\mathbb{Z}/4m\mathbb{Z})^{\times}\longrightarrow 0,$
+where the homomorphism from $\Gamma_{0}(4m)$ to $(\mathbb{Z}/4m\mathbb{Z})^{\times}$ is given by the assignment $\big{(}\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\big{)}\mapsto d$.
+From this we conclude that
+$[\Gamma_{0}:\Gamma_{1}]$ $=[\mathrm{SL}_{2}(\mathbb{Z}):\Gamma_{0}(4m)]\cdot[\Gamma_{0}(4m):\Gamma_{0,1}(4m)]$
+$=[\mathrm{SL}_{2}(\mathbb{Z}):\Gamma_{0}(4m)]\cdot\big{|}(\mathbb{Z}/4m\mathbb{Z})^{\times}\big{|}=[\mathrm{SL}_{2}(\mathbb{Z}):\Gamma_{0}(4m)]\cdot\varphi(m),$
+where $\varphi(\cdot)$ denotes Euler’s $\varphi$-function. Since we trivially have $\varphi(m)\leq m$ and since it is well known that
+$[\mathrm{SL}_{2}(\mathbb{Z}):\Gamma_{0}(4m)]=O_{\epsilon}\big{(}m^{1+\epsilon}\big{)}$
+with an implied constant depending on the choice of $\epsilon>0$, the claim follows. ∎
 Theorem 4.4.
-For $k\in\mathbb{Z}_{\geq 5}$ and $m\in\mathbb{Z}_{\geq 1}$, let $f\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$ be a Jacobi cusp form of weight $k$ and index $m$ for the full modular group $\Gamma_{0}=\mathrm{SL}_{2}(\mathbb{Z})$, which is normalized with respect to the Petersson inner product.
-Then, we have the bound
-(61) $\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|f(\tau,z)\|^{2}_{\mathrm{Pet}}=O_{\epsilon}\big{(}k^{\frac{3}{2}}\,m^{2+\epsilon}\big{)},$
-where the implied constant depends only on the choice of $\epsilon>0$.
+For $k\in\mathbb{Z}_{\geq 5}$ and $m\in\mathbb{Z}_{\geq 1}$, let $\phi\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$ be an $L^{2}$-normalized Jacobi cusp form of weight $k$ and index $m$ for the full modular group $\Gamma_{0}=\mathrm{SL}_{2}(\mathbb{Z})$.
+Then, we have the $L^{\infty}$-norm bound
+(49) $\|\phi\|^{2}_{L^{\infty}}=\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|\phi(\tau,z)\|^{2}_{\mathrm{Pet}}=O_{\Gamma_{0},\epsilon}\big{(}k^{2}\,m^{\frac{7}{2}+\epsilon}\big{)},$
+where the implied constant depends on $\Gamma_{0}$ and the choice of $\epsilon>0$.
 Proof.
-Substituting the decomposition (13) of the Jacobi form $f(\tau,z)$ into its pointwise Petersson norm and applying the Cauchy–Schwartz inequality, we compute
-$\|f(\tau,z)\|^{2}_{\mathrm{Pet}}$ $=\bigg{|}\sum_{\mu=0}^{2m-1}\varphi_{\mu}(\tau)\vartheta_{\mu,m}(\tau,z)\bigg{|}^{2}\,\eta^{k}\,e^{-\frac{4\pi my^{2}}{\eta}}$
-$\leq\bigg{(}\sum_{\mu=0}^{2m-1}|\varphi_{\mu}(\tau)|^{2}\,\eta^{k-1/2}\bigg{)}\bigg{(}\sum_{\mu=0}^{2m-1}|\vartheta_{\mu,m}(\tau,z)|\,\eta^{1/2}\,e^{-\frac{4\pi my^{2}}{\eta}}\bigg{)}$
-$=\bigg{(}\sum_{\mu=0}^{2m-1}\|\varphi_{\mu}(\tau)\|^{2}_{\mathrm{Pet}}\bigg{)}\bigg{(}\sum_{\mu=0}^{2m-1}\|\vartheta_{\mu,m}(\tau,z)\|_{\mathcal{L}_{\tau}^{\otimes{m}}}^{2}\bigg{)}.$
-Thus
-(62) $\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|f(\tau,z)\|^{2}_{\mathrm{Pet}}\leq\sup_{\tau\in\mathbb{H}}\sum_{\mu=0}^{2m-1}\|\varphi_{\mu}(\tau)\|^{2}_{\mathrm{Pet}}\cdot\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\sum_{\mu=0}^{2m-1}\|\vartheta_{\mu,m}(\tau,z)\|_{\mathcal{L}_{\tau}^{\otimes{m}}}^{2}.$
-Since $f$ is normalized with respect to the Petersson inner product, we have from (15) that
-(63) $\langle f,f\rangle_{\mathrm{Pet}}=\frac{1}{\sqrt{4m}\,[\Gamma_{0}:\Gamma_{1}]}\sum_{\mu=0}^{2m-1}\langle\varphi_{\mu},\varphi_{\mu}\rangle_{\mathrm{Pet}}=1.$
-Combining (63) with the bound (35) given in Theorem 3.5, we find the bound
-(64) $\sum_{\mu=0}^{2m-1}\|\varphi_{\mu}(\tau)\|^{2}_{\mathrm{Pet}}=\sum_{\mu=0}^{2m-1}\langle\varphi_{\mu},\varphi_{\mu}\rangle_{\mathrm{Pet}}\,\frac{\|\varphi_{\mu}(\tau)\|^{2}_{\mathrm{Pet}}}{\langle\varphi_{\mu},\varphi_{\mu}\rangle_{\mathrm{Pet}}}\leq C\,\sqrt{4m}\,[\Gamma_{0}:\Gamma_{1}]\,k^{\frac{3}{2}},$
-where $C$ is a positive constant depending on $\Gamma_{0}$. Recalling now that
-(65) $[\Gamma_{0}:\Gamma_{1}]=[\mathrm{SL}_{2}(\mathbb{Z}):\Gamma_{0}(4m)]=O_{\epsilon}\big{(}m^{1+\epsilon}\big{)},$
-where the implied constant depends only on the choice of $\epsilon>0$, we derive from the bounds (64) and (65), the bound
-(66) $\sup_{\tau\in\mathbb{H}}\sum_{\mu=0}^{2m-1}\|\varphi_{\mu}(\tau)\|^{2}_{\mathrm{Pet}}=O_{\epsilon}\big{(}k^{\frac{3}{2}}\,m^{\frac{3}{2}+\epsilon}\big{)},$
-where the implied constant depends only on the choice of $\epsilon>0$.
-The claimed bound (61) finally follows from (62) by combining the bound (66) with the bound (59) established in Proposition 4.3.
-∎
+Substituting the decomposition (13) of the Jacobi form $\phi(\tau,z)$ into its pointwise Petersson norm and applying the Cauchy–Schwartz inequality, we find the estimate
+$\|\phi(\tau,z)\|^{2}_{\mathrm{Pet}}$ $=\bigg{|}\sum_{\mu=0}^{2m-1}h_{\mu}(\tau)\vartheta_{\mu,m}(\tau,z)\bigg{|}^{2}\,\eta^{k}\,e^{-\frac{4\pi my^{2}}{\eta}}$
+$\leq\bigg{(}\sum_{\mu=0}^{2m-1}|h_{\mu}(\tau)|^{2}\,\eta^{k-\frac{1}{2}}\bigg{)}\bigg{(}\sum_{\mu=0}^{2m-1}|\vartheta_{\mu,m}(\tau,z)|^{2}\,\eta^{\frac{1}{2}}\,e^{-\frac{4\pi my^{2}}{\eta}}\bigg{)}$
+$=\bigg{(}\sum_{\mu=0}^{2m-1}\|h_{\mu}(\tau)\|^{2}_{\mathrm{Pet}}\bigg{)}\bigg{(}\sum_{\mu=0}^{2m-1}\|\vartheta_{\mu,m}(\tau,z)\|_{\mathrm{Pet}}^{2}\bigg{)}.$
+Thus, we arrive by means of Lemma 4.1 at the bound
+$\|\phi(\tau,z)\|^{2}_{\mathrm{Pet}}$ $\leq\sum_{\mu=0}^{2m-1}\|h_{\mu}(\tau)\|^{2}_{\mathrm{Pet}}\cdot\sum_{\mu=0}^{2m-1}\|\vartheta_{\mu,m}(\tau,z)\|_{\mathrm{Pet}}^{2}$
+$\leq 2m\sum_{\mu=0}^{2m-1}\|h_{\mu}(\tau)\|^{2}_{\mathrm{Pet}}\,\eta^{\frac{1}{2}}+O\bigg{(}m^{\frac{1}{2}}\,\sum_{\mu=0}^{2m-1}\|h_{\mu}(\tau)\|^{2}_{\mathrm{Pet}}\bigg{)},$
+from which we derive, after recalling the comment at the beginning of this section,
+$\|\phi\|^{2}_{L^{\infty}}$ $=\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|\phi(\tau,z)\|^{2}_{\mathrm{Pet}}=\sup_{\begin{subarray}{c}\tau\in\mathcal{F}_{\Gamma_{0}}\\ z\in E_{\tau}\end{subarray}}\|\phi(\tau,z)\|^{2}_{\mathrm{Pet}}$
+(50) $\leq 2m\sum_{\mu=0}^{2m-1}\sup_{\tau\in\mathcal{F}_{\Gamma_{0}}}\big{(}\|h_{\mu}(\tau)\|^{2}_{\mathrm{Pet}}\,\eta^{\frac{1}{2}}\big{)}+O\bigg{(}m^{\frac{1}{2}}\sum_{\mu=0}^{2m-1}\sup_{\tau\in\mathcal{F}_{\Gamma_{0}}}\|h_{\mu}(\tau)\|^{2}_{\mathrm{Pet}}\bigg{)}.$
+In order to be able to apply Lemma 4.2 and Theorem 3.5 to the two summands in (50), respectively, we need to $L^{2}$-normalize the modular forms $h_{\mu}$ ($\mu=0,\ldots,2m-1$) under consideration.
+To do so, we observe that formula (15) in conjunction with the $L^{2}$-normalization of $\phi$ gives
+(51) $\|\phi\|^{2}_{L^{2}}=\frac{1}{\sqrt{4m}\,[\Gamma_{0}:\Gamma_{1}]}\sum_{\mu=0}^{2m-1}\|h_{\mu}\|^{2}_{L^{2}}=1.$
+An application of Lemma 4.3 to (51) then gives
+$\sum_{\mu=0}^{2m-1}\|h_{\mu}\|^{2}_{L^{2}}=\sqrt{4m}\,[\Gamma_{0}:\Gamma_{1}]=O_{\epsilon}\big{(}m^{\frac{5}{2}+\epsilon}\big{)}$
+with an implied constant depending on the choice of $\epsilon>0$. With regard to the first summand in (50), we thus obtain by Lemma 4.2 the bound
+$2m\sum_{\mu=0}^{2m-1}\sup_{\tau\in\mathcal{F}_{\Gamma_{0}}}\big{(}\|h_{\mu}(\tau)\|^{2}_{\mathrm{Pet}}\,\eta^{\frac{1}{2}}\big{)}$ $=2m\sum_{\mu=0}^{2m-1}\sup_{\tau\in\mathcal{F}_{\Gamma_{0}}}\bigg{(}\frac{\|h_{\mu}(\tau)\|^{2}_{\mathrm{Pet}}}{\|h_{\mu}\|^{2}_{L^{2}}}\,\eta^{\frac{1}{2}}\bigg{)}\|h_{\mu}\|^{2}_{L^{2}}$
+(52) $=O_{\Gamma_{0}}\big{(}2mk^{2}\big{)}\sum_{\mu=0}^{2m-1}\|h_{\mu}\|^{2}_{L^{2}}=O_{\Gamma_{0},\epsilon}\big{(}k^{2}\,m^{\frac{7}{2}+\epsilon}\big{)}$
+with an implied constant depending on $\Gamma_{0}$ and the choice of $\epsilon>0$. With regard to the second summand in (50), we derive from Theorem 3.5 the bound
+$m^{\frac{1}{2}}\sum_{\mu=0}^{2m-1}\sup_{\tau\in\mathcal{F}_{\Gamma_{0}}}\|h_{\mu}(\tau)\|^{2}_{\mathrm{Pet}}$ $=m^{\frac{1}{2}}\sum_{\mu=0}^{2m-1}\sup_{\tau\in\mathcal{F}_{\Gamma_{0}}}\bigg{(}\frac{\|h_{\mu}(\tau)\|^{2}_{\mathrm{Pet}}}{\|h_{\mu}\|^{2}_{L^{2}}}\bigg{)}\|h_{\mu}\|^{2}_{L^{2}}$
+(53) $=O_{\Gamma_{0}}\big{(}m^{\frac{1}{2}}k^{\frac{3}{2}}\big{)}\sum_{\mu=0}^{2m-1}\|h_{\mu}\|^{2}_{L^{2}}=O_{\Gamma_{0},\epsilon}\big{(}k^{\frac{3}{2}}\,m^{3+\epsilon}\big{)}$
+with an implied constant depending on $\Gamma_{0}$ and the choice of $\epsilon>0$. The bounds (52) and (53) complete the proof of the theorem. ∎
 Remark 4.5.
-The bound (61) is polynomial in $k$ and $m$, and thus improves W.
-Kohnen’s bound (1), which is exponential in $k$.
 Comparing our bound with the one obtained by P.
 Anamby and S.
-Das in [AD23], there is an improvement with regard to the polynomial growth in $k$, while the polynomial growth in $m$ is slightly worse.
+Das in [AD23], we obtain the same polynomial growth in $k$, while the polynomial growth of their bound is better by a factor of $m^{\frac{3}{4}}$.
 Acknowledgements
@@
 Lastly, all the three authors would like to acknowledge the hospitality of the International Centre for Theoretical Sciences (ICTS) during a visit for participating in the program Algebraic and Analytic Aspects of Automorphic Forms (Code: ICTS/aforms2019/02), where this project was initiated.
-Finally, the authors express their gratitude to Edgar Assing, whose observations have helped to improve our initial bound (3) by a factor of $\sqrt{m}$.